### v1

# and Qian Liu

Abstract
Reconfigurable antennas possess the capability to dynamically adjust their fundamental operating characteristics, thereby enhancing system adaptability and performance. To fully exploit this flexibility in modern wireless communication systems, this paper considers a novel tri-hybrid beamforming architecture, which seamlessly integrates pattern-reconfigurable antennas with both analog and digital beamforming. The proposed tri-hybrid architecture operates across three layers: (i) a radiation beamformer in the electromagnetic (EM) domain for dynamic pattern alignment, (ii) an analog beamformer in the radio-frequency (RF) domain for array gain enhancement, and (iii) a digital beamformer in the baseband (BB) domain for multi-user interference mitigation. To establish a solid theoretical foundation, we first develop a comprehensive mathematical model for the tri-hybrid beamforming system and formulate the signal model for a multi-user multi-input single-output (MU-MISO) scenario. The optimization objective is to maximize the sum-rate while satisfying practical constraints. Given the challenges posed by high pilot overhead and computational complexity, we introduce an innovative tri-timescale beamforming framework, wherein the radiation beamformer is optimized over a long-timescale, the analog beamformer over a medium-timescale, and the digital beamformer over a short-timescale. This hierarchical strategy effectively balances performance and implementation feasibility. Simulation results validate the performance gains of the proposed tri-hybrid architecture and demonstrate that the tri-timescale design significantly reduces pilot overhead and computational complexity, highlighting its potential for future wireless communication systems.
Index Terms:
Reconfigurable antennas, tri-hybrid beamforming, tri-timescale, hybrid beamforming.
I Introduction
Future wireless communication systems are expected to leverage ultra-wide spectrum resources in high-frequency bands, such as millimeter wave (mmWave), to accommodate the growing demand for ultra-high data rates [1]. To mitigate the severe propagation loss in these frequency bands, massive multiple-input multiple-output (mMIMO) systems or extremely large antenna arrays (ELAA) are typically employed to ensure reliable signal transmission [2]. In conventional antenna arrays, each antenna element typically exhibits a fixed and identical radiation patterns that concentrate energy towards predetermined directions, often perpendicular to the array. Within this framework, most existing studies focus on signal processing-based beamforming techniques, which adjust the phase and amplitude of signals across antenna elements to direct beam towards specific directions.
However, the limitations of conventional pattern-fixed antennas are becoming increasingly apparent in emerging applications, such as ELAA and near-field communications [3]. Specifically, system performance is significantly influenced not only by array beamforming but also by the radiation gain of each individual antenna element. While antennas with narrow, highly energy-concentrated radiation patterns provide high gain within specific angular ranges, they suffer substantial performance degradation in the endfire direction due to a sharp decrease in signal strength. Conversely, antennas with broader radiation patterns provide wider coverage at the cost of reduced gain, ultimately compromising the quality of service.
Reconfigurable antenna technology offers a promising solution to the challenges faced by conventional pattern-fixed antennas by enhancing adaptability [4]-[7]. Reconfigurable antennas can adapt their internal parameters to modify key operating characteristics such as radiation patterns [8]-[10], frequency responses [11], and polarizations [12], [13]. Among these, radiation pattern reconfiguration is particularly crucial for wireless communication systems, as it directly impacts both the transmitted and received signal strength. To enable flexible control over these characteristics, several hardware implementations have been developed, such as electronically steerable parasitic array radiator (ESPAR) antennas [14], frequency selective surfaces (FSS) [15], and pixel antennas [16], all of which share similar underlying principles and architectures. Typically, a reconfigurable antenna consists of an active element and multiple parasitic elements, where the connection states of parasitic elements are dynamically controlled to alter the current distribution within the antenna, thereby modifying its electromagnetic (EM) properties. This capability introduces new degrees of freedom (DoF) in the EM domain, offering the potential to overcome existing performance limitations and enable more efficient wireless communication systems.
With significant advancements in reconfigurable antenna technology, integrating pattern-reconfigurable antennas into antenna arrays has gained considerable attention in recent years. For instance, the authors in [17] employ multi-port network theory and beamspace channel representation to develop an EM-based communication model for pixel antennas, which is then used to optimize antenna coding for maximizing channel gain. A similar approach is applied in [14] to construct an ESPAR antenna model, enabling analysis and optimization of ESPAR beamforming. Additionally, the authors in [18] introduce a decomposition method based on spherical harmonic functions to model radiation patterns in the EM domain. This approach decomposes the radiation pattern into a linear combination of orthogonal bases, facilitating the iterative design of radiation patterns and digital precoders. While these studies highlight the advantages of replacing conventional pattern-fixed antennas with innovative pattern-reconfigurable ones, they overlook the significant hardware costs and power consumption incurred by the increasing number of antennas and operating frequencies. To tackle these issues, existing literature commonly suggests hybrid beamforming architectures for mMIMO systems, as a power-efficient alternative to conventional fully-digital beamforming [19].
In hybrid beamforming, analog beamforming controls the beam direction in the radio frequency (RF) domain, while digital beamforming performs spatial signal processing in the baseband (BB) domain, leveraging a reduced number of RF chains and hardware-efficient phase shifters (PSs). Common hybrid beamforming architectures, including fully-connected [20], partially-connected [21], and dynamic architectures [22], strike a balance between performance and implementation cost. While recent advances in hardware and power-efficient hybrid architectures have enabled large-scale MIMO deployments for 6G and beyond, the limitations of conventional pattern-fixed antennas remain a significant challenge. To address this issue, we propose leveraging the flexibility of pattern-reconfigurable antenna technology to address the performance degradation caused by fixed radiation patterns, while utilizing hybrid beamforming to reduce power consumption without sacrificing performance.
In this paper, we consider a novel tri-hybrid beamforming architecture that integrates pattern-reconfigurable antennas with both analog and digital beamforming. Specifically, this tri-hybrid architecture consists of three layers: A radiation beamformer in the EM domain to align radiation patterns with desired direction, an analog beamformer in the RF domain to enhance array gain, and a digital beamformer in the BB domain to mitigate multi-user interference. The concept of tri-hybrid architecture was first introduced in [23], [24]. While prior analyses and simulations have demonstrated the advantages of tri-hybrid beamforming in terms of spectral efficiency and energy consumption, several critical issues must be addressed before its practical implementation in future wireless systems. First, accurate acquisition of instantaneous channel state information (CSI) becomes more challenging due to the higher dimensionality of CSI and the limited number of RF chains, leading to increased pilot overhead and requiring advanced estimation algorithms. Second, optimizing the beamformers across all three domains is more complicated due to parameter coupling, discrete phase-shift constraints, and the nonlinear characteristics of reconfigurable antennas, etc.
Recently, a two-timescale framework has been proposed to address system overhead and design complexity in conventional hybrid beamforming. In this framework, the analog beamformer is optimized based on channel statistics over a long-timescale, while the digital beamformer is designed using low-dimensional instantaneous effective CSI over a short-timescale [25]-[27]. Building on this concept and leveraging the unique characteristics of the tri-hybrid architecture, we propose a novel tri-timescale framework for tri-hybrid beamforming design, effectively managing practical overhead and complexity. The main contributions of this paper are summarized as follows:
  • We first develop a comprehensive mathematical model for the radiation beamformer based on multiple pattern-reconfigurable antennas, along with the corresponding virtual angular index channel in the EM domain. Building upon this model, we derive the formulation of the tri-hybrid beamforming architecture, establish the signal model for a multi-user multiple-input single-output (MU-MISO) system employing this architecture, and present the associated optimization problem.
  • For the tri-hybrid beamforming design, our objective is to maximize the sum-rate while satisfying constraints on radiation pattern energy, constant modulus and discrete phase shifts of PSs, a non-overlapping dynamic connection network, and the total transmit power budget. To effectively address the challenges of high pilot overhead and substantial computational complexity, we propose a novel tri-timescale beamforming framework, where the radiation beamformer is optimized over a long-timescale, the analog beamformer over a medium-timescale, and the digital beamformer over a short-timescale.
  • Extensive simulations are conducted to validate the capability of the tri-hybrid beamforming architecture in enhancing system performance, as well as the effectiveness of the proposed tri-timescale beamforming design algorithm in reducing pilot overhead and computational complexity.
Figure 1: The hardware prototypes of pattern-reconfigurable antenna (left [8], right [9]).
Notations: Boldface lower-case and upper-case letters indicate column vectors and matrices, respectively. $(\cdot)^{*}$, $(\cdot)^{T}$, and $(\cdot)^{H}$ denote the conjugate, transpose, and transpose-conjugate. $|a|$, $\|\mathbf{a}\|$, and $\|\mathbf{A}\|_{F}$ are the magnitude of scalar $a$, the norm of vector $\mathbf{a}$, and the Frobenius norm of matrix $\mathbf{A}$. $\mathrm{vec}(\mathbf{A})$ stacks $\mathbf{A}$’s columns into a long column vector. Notation $\otimes$ is the Kronecker product of matrices. $\mathrm{blkdiag}\{\mathbf{A}_{1},\mathbf{A}_{2},\ldots,\mathbf{A}_{M}\}$ denotes a block diagonal matrix. $\mathbf{1}_{M}$ and $\mathbf{I}_{M}$ represent an $M\!\times\!1$ vector of ones and an $M\!\times\!M$ identity matrix, respectively. ${}\Re\{\cdot\}$ denotes the real part of a complex number. The statistical expectation is given by $\mathbb{E}\{\cdot\}$. $\mathbf{a}(i)$ denotes the $i$-th entry of the vector $\mathbf{a}$. $\mathbf{A}(i,:)$ and $\mathbf{A}(i,j)$ denote the $i$-th row and $(i,j)$-th element of matrix $\mathbf{A}$, respectively.
II System Model and Problem Formulation
Figure 2: The illustration of a pattern-reconfigurable antenna.
II-A Pattern-reconfigurable Antenna
Pattern-reconfigurable antenna is an advanced technology that enables dynamic adaptation of radiation patterns to enhance signal quality, mitigate interference, and improve coverage in wireless communication systems. Various hardware implementations have been explored to achieve flexible radiation pattern control, including ESPAR antennas [14], FSS [15], and pixel antennas [16]. Among these, pixel antenna technology has emerged as the most practical and versatile solution for integration into compact antenna arrays. The hardware prototypes of pixel-based reconfigurable antennas are illustrated in Fig. 1.
Having demonstrated their feasibility in real-world applications, we now describe the hardware configuration and operational principle of the pixel-based pattern-reconfigurable antennas. As illustrated in Fig. 2, pixel antenna technology discretizes a continuous antenna surface into electrically small parasitic elements, called pixels, with a central radiating source driving the entire antenna. These pixels serving as fundamental building blocks, utilize low-voltage, binary-controlled pin diodes to connect or disconnect adjacent pixels. By adjusting the states of these switches, different pixel topologies are formed, which modify the current distribution across the antenna surface. This reconfiguration alters the electric and magnetic field components radiated by the antenna, enabling the generation of a diverse set of radiation patterns.
To evaluate the performance improvements brought by pattern-reconfigurable antennas, it is crucial to establish a mathematical model for both the radiation pattern and the corresponding channel characteristics. As aforementioned, pattern-reconfigurable antennas introduce an additional DoF in the EM domain, effectively expanding the information coverage dimension of conventional antennas, from a one-dimensional spatial representation to an $M$-dimensional angular information, where $M$ represents the number of uniform sampling points in the angular domain. Specifically, a single reconfigurable radiation pattern (exemplified by the red pattern in Fig. 2) is characterized over $M$ discrete directions in the EM domain and is represented by $\mathbf{f}_{\mathrm{EM}}\in\mathbb{R}^{M}$, with each element corresponding to the radiation gain in a specific direction. To model reception characteristics, we introduce the virtual angular index channel for the pattern-reconfigurable antenna, denoted as $\mathbf{h}_{\mathrm{EM}}\in\{0,1\}^{M}$, which serves as an index matrix indicating the spatial angles at which the user is located. Consequently, the effective gain in the EM domain can be expressed as $G_{\mathrm{EM}}=\mathbf{h}_{\mathrm{EM}}^{H}\mathbf{f}_{\mathrm{EM}}$, which quantifies the radiation intensity directed toward the user’s actual spatial angle. By optimizing the radiation pattern $\mathbf{f}_{\mathrm{EM}}$ to align with the angular domain channel characteristics, the correlation between them can be maximized, leading to enhanced radiation efficiency. This formulation serves as the foundation for the following theoretical analysis and optimization.
Figure 3: Tri-hybrid beamforming architecture.
Inspired by the increased design DoF introduced by pattern-reconfigurable antennas, we integrate this advanced hardware component into the conventional hybrid beamforming architecture and propose a tri-hybrid beamforming framework that incorporates beamforming across three domains. Specifically, radiation beamforming in the EM domain is employed to dynamically shape the radiation pattern, analog beamforming in the RF domain is utilized to enhance array gain, and digital beamforming in the BB domain is implemented to mitigate multi-user interference.
II-B Tri-hybrid Beamforming Architecture
To provide a clear illustration of the proposed tri-hybrid beamforming architecture, we consider a mmWave MU-MISO downlink communication system, as shown in Fig. 3. Specifically, a base station (BS) equipped with a uniform linear array (ULA) serves multiple single-antenna user equipments (UEs). The BS employs reconfigurable antennas that dynamically adjust their radiation patterns, while each UE is equipped with a conventional omnidirectional antenna. The antenna spacing at the BS is set to $d={\lambda}/{2}$, where $\lambda$ is the center carrier wavelength. Let $\mathcal{N}_{\mathrm{t}}\triangleq\{1,2,\ldots,N_{\mathrm{t}}\}$ denote the set of BS antennas, $\mathcal{K}\triangleq\{1,2,\ldots,K\}$ denote the set of UEs, and $\mathcal{L}\triangleq\{1,2,\ldots,L\}$ denote the set of propagation paths between the BS and each UE. In the following, we mathematically formulate the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in the EM domain, the analog beamformer $\mathbf{F}_{\mathrm{RF}}$ in the RF domain and the digital beamformer $\mathbf{F}_{\mathrm{BB}}$ in the BB domain, establishing the foundation for the tri-hybrid beamforming architecture.
EM domain: Building upon the radiation pattern model of a single pattern-reconfigurable antenna $\mathbf{f}_{\mathrm{EM}}$ presented in Section II-A, we extend this model to characterize the programmable radiation pattern of the entire antenna array at the BS, which consists of $N_{\mathrm{t}}$ reconfigurable antennas. The resulting radiation beamformer in the EM domain is expressed in the following matrix form
$\mathbf{F}_{\mathrm{EM}}\triangleq\mathrm{blkdiag}\{\mathbf{f}_{\mathrm{EM},1},\mathbf{f}_{\mathrm{EM},2},\ldots,\mathbf{f}_{\mathrm{EM},N_{\mathrm{t}}}\}\in\mathbb{R}^{MN_{\mathrm{t}}\times N_{\mathrm{t}}},$ (1)
where the $m$-th element of $\mathbf{f}_{\mathrm{EM},n}$ represents the radiation gain of the $n$-th antenna in the $m$-th angular direction. To ensure fairness in pattern design, the total energy of each antenna’s radiation pattern is restricted and normalized as $\|\mathbf{f}_{\mathrm{EM},n}\|^{2}=1,\forall n$. To further characterize the spatial mapping of the reconfigurable antennas, we introduce the virtual angular index channel matrix
$\overline{\mathbf{H}}_{\mathrm{EM},k}\triangleq\mathbf{I}_{N_{\mathrm{t}}}\otimes\mathbf{H}_{\mathrm{EM},k}\in\{0,1\}^{MN_{\mathrm{t}}\times LN_{\mathrm{t}}},~{}\forall k,$ (2)
where $\mathbf{H}_{\mathrm{EM},k}\triangleq[\mathbf{h}_{\mathrm{EM},k,1},\mathbf{h}_{\mathrm{EM},k,2},\ldots,\mathbf{h}_{\mathrm{EM},k,L}]\in\{0,1\}^{M\times L}$ that specifies the spatial angles associated with the $k$-th UE. Each column $\mathbf{h}_{\mathrm{EM},k,l}$ contains exactly one non-zero element, indicating the angular direction of the $l$-th propagation path for the $k$-th UE. Consequently, the radiation gain $\mathbf{G}_{\mathrm{EM},k}\in\mathbb{R}^{LN_{\mathrm{t}}\times N_{\mathrm{t}}}$ of all antennas towards the $k$-th UE can be formulated as
$\mathbf{G}_{\mathrm{EM},k}$ $=\overline{\mathbf{H}}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}$ (3)
$=\mathrm{blkdiag}\{\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{f}_{\mathrm{EM},1},\ldots,\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{f}_{\mathrm{EM},N_{\mathrm{t}}}\},$
where the $l$-th entry of $\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{f}_{\mathrm{EM},n}$ represents the actual radiation gain of the $n$-th antenna towards the $l$-th direction for the $k$-th UE. In summary, the matrices $\mathbf{F}_{\mathrm{EM}}$ and $\overline{\mathbf{H}}_{\mathrm{EM},k}$ mathematically capture the additional degrees of freedom introduced by reconfigurable antennas, enabling a more flexible and adaptive radiation beamforming strategy in the EM domain.
RF domain: We adopt a dynamic-subarray analog beamforming architecture, where each RF chain is adaptively connected to a disjoint subset of the total transmit antennas via a switch network. By fully exploiting the design flexibility of dynamic antenna selection, this architecture achieves an optimal balance between system efficiency and communication performance. Without loss of generality, we assume that the number of RF chains $N_{\mathrm{RF}}$ is not less than the number of UEs $K$, i.e., $N_{\mathrm{RF}}\geq K$, and the set of RF chains is denoted as $\mathcal{N}_{\mathrm{RF}}\triangleq\{1,2,\ldots,N_{\mathrm{RF}}\}$. The analog beamformer employs a set of low-resolution PSs, each following a constant modulus constraint with quantized phase control of $B$ bits. The discrete phase set of PSs is given by $\mathcal{F}\triangleq\{\frac{1}{\sqrt{N_{\mathrm{t}}}}e^{\jmath\frac{2\pi b}{2^{B}}}|b=0,1,\ldots,2^{B}-1\}$. To facilitate dynamic-subarray beamforming, we define the analog beamformer matrix as $\mathbf{F}_{\mathrm{RF}}\in\{\mathcal{F},0\}^{N_{\mathrm{t}}\times N_{\mathrm{RF}}}$, which implies that if the $l_{\mathrm{RF}}$-th RF chain is connected to the $n$-th antenna, the corresponding phase-shift takes a non-zero value, i.e., $\mathbf{F}_{\mathrm{RF}}(n,l_{\mathrm{RF}})\in\mathcal{F}$; if not, it is set as zero, i.e., $\mathbf{F}_{\mathrm{RF}}(n,l_{\mathrm{RF}})=0$. Moreover, to ensure non-overlapping subarrays, each antenna is dynamically assigned to only one RF chain, meaning that the analog beamformer matrix is restricted to have only one non-zero element per row, i.e., $\|\mathbf{F}_{\mathrm{RF}}(n,:)\|_{0}=1$.
BB domain: We exploit the digital beamformer $\mathbf{F}_{\mathrm{BB}}\triangleq[\mathbf{f}_{\mathrm{BB},1},\mathbf{f}_{\mathrm{BB},2},\ldots,\mathbf{f}_{\mathrm{BB},K}]\in\mathbb{C}^{N_{\mathrm{RF}}\times K}$ to pre-process transmitted signals in the BB domain, addressing the limitation of having fewer RF chains, which restricts direct control over each antenna’s signal. This procedure effectively mitigates multi-user interference and enhances overall system performance.
II-C Channel Model
After establishing the mathematical formulation of the tri-hybrid beamforming architecture, our focus now shifts to developing a comprehensive channel model that accommodates the expanded dimensionality introduced by the pattern-reconfigurable antenna array. Specifically, unlike conventional channel models that primarily account for path loss and spatial phase gradients, the proposed model incorporates the virtual angular index channel matrix $\overline{\mathbf{H}}_{\mathrm{EM},k},\forall k$ in (2). This integration enables an explicit representation of the angular characteristics of each propagation path. Furthermore, when combined with the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in (1), it enables the determination of the radiation intensity $\mathbf{G}_{\mathrm{EM},k},\forall k$ in (3), facilitating a more precise characterization of the system’s electromagnetic behavior.
To establish a foundation for the innovative channel model of the pattern-reconfigurable antenna array, we first present the traditional mmWave channel model. Due to the sparse scattering characteristics of the mmWave band, the channel typically consists of a few dominant clusters, each containing multiple propagation paths [28]. Let $\mathcal{C}\triangleq\{1,2,\ldots,C\}$ denote the set of scattering clusters, and $\mathcal{L}_{c}\triangleq\{1,2,\ldots,L_{c}\}$ denote the set of multi-path rays within the $c$-th scattering cluster. The channel between the $n$-th antenna and the $k$-th UE can be expressed as
$h_{k,n}=\sum_{c=1}^{C}\sum_{l_{c}=1}^{L_{c}}\alpha_{k,c,l_{c}}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,c,l_{c}}},~{}\forall k,n,$ (4)
where $\theta_{k,c,l_{c}}$ and $\alpha_{k,c,l_{c}}$ denote the angle of departure (AoD) and channel gain corresponding to the $l_{c}$-th ray in the $c$-th scattering cluster, respectively. The AoD $\theta_{k,c,l_{c}}$ can be described as
$\theta_{k,c,l_{c}}=\overline{\theta}_{k,c}+\Delta\theta_{k,l_{c}},~{}\forall k,c,l_{c},$ (5)
where $\overline{\theta}_{k,c}$ denotes the nominal AoD of the $c$-th scattering cluster for the $k$-th UE, and $\Delta\theta_{k,l_{c}}$ denotes the angular deviation of the $l_{c}$-th ray from the nominal direction, following Gaussian distribution $\Delta\theta_{k,l_{c}}\sim\mathcal{N}(0,\varsigma^{2}_{k,c})$, with $\varsigma_{k,c}$ indicating the angle spread. The channel gain involves the typically distance-dependent path loss with $r_{k,c,l_{c}}$ representing the propagation distance. It follows the same modeling as in (5), i.e., $r_{k,c,l_{c}}=\overline{r}_{k,c}+\Delta r_{k,l_{c}}$, where $\Delta r_{k,l_{c}}\sim\mathcal{N}(0,\iota^{2}_{k,c})$. For better clarity in the subsequent channel model formulation, we define $L\triangleq\sum_{c=1}^{C}L_{c}$ as the total number of paths. With this notation, the channel model in (4) can be reformulated in a more compact form as follows
$h_{k,n}=\sum_{l=1}^{L}\alpha_{k,l}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,l}},~{}\forall k,n.$ (6)
Furthermore, the integration of pattern-reconfigurable antennas extends the channel representation from a one-dimensional to an $M$-dimensional formulation by incorporating additional EM domain channel information, i.e., $\overline{\mathbf{H}}_{\mathrm{EM},k},\forall k$ in (2). Specifically, the dimension-extended channel for the $n$-th pattern-reconfigurable antenna, denoted as $\mathbf{h}_{k,n}\in\mathbb{C}^{M}$, is expressed as follows
$\mathbf{h}_{k,n}=\sum_{l=1}^{L}(\alpha_{k,l}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,l}})\otimes\mathbf{h}_{\mathrm{EM},k,l},~{}\forall k,n.$ (7)
Following this formulation, by stacking the channel vectors $\mathbf{h}_{k,n},\forall n$ of all $N_{\mathrm{t}}$ antennas, we obtain the complete channel vector for the $k$-th UE, given by $\mathbf{h}_{k}\triangleq[\mathbf{h}_{k,1}^{T},\mathbf{h}_{k,2}^{T},\ldots,\mathbf{h}_{k,N_{\mathrm{t}}}^{T}]^{T}\in\mathbb{C}^{MN_{\mathrm{t}}}$. The corresponding expression can be explicitly formulated as
$\mathbf{h}_{k}=\sum_{l=1}^{L}(\alpha_{k,l}\mathbf{a}_{k,l})\otimes\mathbf{h}_{\mathrm{EM},k,l},~{}\forall k,$ (8)
where
$\mathbf{a}_{k,l}$ $\triangleq[1,e^{-\jmath\frac{2\pi}{\lambda}d\sin\theta_{k,l}},\ldots,e^{-\jmath\frac{2\pi}{\lambda}(N_{\mathrm{t}}-1)d\sin\theta_{k,l}}]^{T},~{}\forall k,l,$ (9)
denotes the steering vector. By applying the Kronecker product formula and rearranging the order of summation, (8) can be equivalently expressed as
$\mathbf{h}_{k}=\mathrm{vec}(\mathbf{H}_{\mathrm{EM},k}\mathbf{H}_{\mathrm{S},k}^{T}),~{}\forall k,$ (10)
with
$\mathbf{H}_{\mathrm{S},k}$ $=[\alpha_{k,1}\mathbf{a}_{k,1},\alpha_{k,2}\mathbf{a}_{k,2},\ldots,\alpha_{k,L}\mathbf{a}_{k,L}]\in\mathbb{C}^{N_{\mathrm{t}}\times L},~{}\forall k,$ (11)
representing the spatial domain channel. Additionally, to enhance clarity and brevity, the EM domain channel and spatial domain channel can be separated into two independent multiplicative expressions, which yields
$\mathbf{h}_{k}=\overline{\mathbf{H}}_{\mathrm{EM},k}\overline{\mathbf{h}}_{\mathrm{S},k},~{}\forall k,$ (12)
where $\overline{\mathbf{H}}_{\mathrm{EM},k}\in\{0,1\}^{MN_{\mathrm{t}}\times LN_{\mathrm{t}}}$ is defined in (2) containing the EM domain channel information, and $\overline{\mathbf{h}}_{\mathrm{S},k}\triangleq\mathrm{vec}(\mathbf{H}_{\mathrm{S},k}^{T})\in\mathbb{C}^{LN_{\mathrm{t}}}$ includes the spatial domain channel information.
II-D System Model
In the MU-MISO downlink system shown in Fig. 3, the transmitted signal is first processed by the digital beamformer $\mathbf{F}_{\mathrm{BB}}$ in the BB domain, followed by the analog beamformer $\mathbf{F}_{\mathrm{RF}}$ in the RF domain and subsequently shaped by the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in the EM domain, before being radiated by the pattern-reconfigurable antennas. After propagating through the spatial channel, it is finally received by the UE. Based on the above mathematical definition, the signal received by the $k$-th UE is given by
$y_{k}$ $=\overline{\mathbf{h}}_{\mathrm{S},k}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}\mathbf{x}+n_{k},~{}\forall k,$ (13)
where $\mathbf{x}\triangleq[x_{1},x_{2},\ldots,x_{K}]^{T}\in\mathbb{C}^{K}$ denotes the transmitted symbols of $K$ UEs that satisfies $\mathbb{E}\{\mathbf{x}\mathbf{x}^{H}\}=\mathbf{I}_{K}$, and $n_{k}\!\sim\!\mathcal{CN}(0,\sigma^{2}_{k})$ denotes the additive white Gaussian noise (AWGN) at the $k$-th UE. The signal-to-interference-plus-noise ratio (SINR) of the $k$-th UE can be calculated as
$\mathrm{SINR}_{k}=\frac{|\mathbf{\overline{h}}_{\mathrm{S},k}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},k}|^{2}}{\sum_{j=1,j\neq k}^{K}|\mathbf{\overline{h}}_{\mathrm{S},k}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},j}|^{2}+\sigma_{k}^{2}},~{}\forall k,\$ (14)
and the corresponding achievable rate can be calculated as
$R_{k}=\log_{2}(1+\mathrm{SINR}_{k}).$ (15)
II-E Problem Formulation
In this paper, based on the tri-hybrid beamforming architecture, we propose to jointly design the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in the EM domain, the analog beamformer $\mathbf{F}_{\mathrm{RF}}$ in the RF domain, and the digital beamformer $\mathbf{F}_{\mathrm{BB}}$ in the BB domain to maximize the sum-rate, while satisfying the constraints of radiation pattern energy, constant modulus and discrete phases of PSs, non-overlapping dynamic connection network, and total transmit power budget. Thus, the joint three domain beamformer optimization problem with full instantaneous CSI is formulated as
$\max_{\mathbf{F}_{\mathrm{EM}},\mathbf{F}_{\mathrm{RF}},\mathbf{F}_{\mathrm{BB}}}$ $\ \ \sum^{K}_{k=1}R_{k}$ (16a)
$\mathrm{s.t.}$ $\ \ \|\mathbf{f}_{\mathrm{EM},n}\|^{2}=1,~{}\forall n,$ (16b)
$\ \ \mathbf{F}_{\mathrm{RF}}(n,l_{\mathrm{RF}})\in\{\mathcal{F},0\},~{}\forall n,l_{\mathrm{RF}},$ (16c)
$\ \ \|\mathbf{F}_{\mathrm{RF}}(n,:)\|_{0}=1,~{}\forall n,$ (16d)
$\ \ \|\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}\|_{F}^{2}\leq P_{\mathrm{t}},$ (16e)
where $P_{\mathrm{t}}$ is the transmit power budget.
Within this tri-hybrid beamforming architecture, accurately acquiring instantaneous CSI is a significant challenge. The analog beamformer in the RF domain suffers from inherent undersampling, while the need to estimate instantaneous CSI for all possible radiation patterns in the EM domain leads to excessive pilot overhead. Moreover, the real-time computation of beamforming matrices across all three domains imposes high computational complexity, severely impacting system processing efficiency. To address these issues, we propose an innovative tri-timescale framework tailored for tri-hybrid beamforming design. This framework decomposes the original non-convex optimization problem in (16) into three independent subproblems, each solvable at a different timescale. This decomposition significantly reduces pilot overhead, enhances computational tractability, and improves overall system efficiency.
III Tri-timescale Beamforming Designs
III-A Tri-timescale Framework
The tri-hybrid beamforming architecture presents significant practical challenges, including high channel estimation overhead and computational complexity. Estimating the full CSI and jointly optimizing beamformers across all three domains in every time-slot is computationally infeasible. To address this, we propose a tri-timescale beamforming framework that effectively balances practical feasibility and system performance, as illustrated in Fig. 4.
The overall time domain corresponds to the channel statistics coherence time, during which the statistical distribution of the channel remains stable. This period is structured hierarchically, with beamforming operations executed across three distinct timescales, as detailed below:
  • Long-timescale: The total duration is divided into $T_{\mathrm{L}}$ super-frames, denoted as $\mathcal{T}_{\mathrm{L}}\triangleq\{1,2,\ldots,T_{\mathrm{L}}\}$. In each super-frame, the BS extracts high-dimensional channel information of the scattering cluster core to design the radiation beamformer, ensuring radiation pattern alignment in the EM domain.
  • Medium-timescale: Each super-frame consists of $T_{\mathrm{M}}$ frames, defined as $\mathcal{T}_{\mathrm{M}}\triangleq\{1,2,\ldots,T_{\mathrm{M}}\}$. In each frame, the BS acquires a statistical channel sample of multi-path rays within the scattering clusters to update the analog beamformer, enhancing array gain in the RF domain.
  • Short-timescale: Each frame is further divided into $T_{\mathrm{S}}$ time-slots, referred to as the channel coherence time, represented as $\mathcal{T}_{\mathrm{S}}\triangleq\{1,2,\ldots,T_{\mathrm{S}}\}$. In each time-slot, the BS obtains an instantaneous low-dimensional effective CSI to refine the digital beamformer, mitigating multi-user interference and accounting for fast channel fluctuations in the BB domain.
This hierarchical framework significantly reduces channel estimation overhead while ensuring efficient beamforming adaptation across different domains and timescales. In the following subsections, we provide a detailed description of the channel estimation processes and beamforming design procedures based on this tri-timescale framework.
Figure 4: The frame structure of tri-timescale tri-hybrid beamforming.
III-B Long-timescale EM Domain Design
III-B.1 Long-timescale EM domain Channel Estimation
Given the specific scattering cluster characteristics described in (4), the angles of the multi-path rays within each scattering cluster exhibit only limited fluctuations around the cluster core. This observation implies that real-time radiation patterns adaptation based on instantaneous CSI is unnecessary. Instead, aligning radiation patterns with the nominal angles of scattering clusters is sufficient to maintain robust communication performance, while significantly reducing channel estimation overhead and beamforming design complexity. Therefore, over the long-timescale, our objective is to estimate the nominal angles and average channel gains of the scattering clusters, which serve as the foundation for radiation beamforming optimization.
Specifically, we employ the Capon algorithm to construct the angular power spectrum (APS) based on the channel covariance matrix, enabling precise spatial domain analysis [29]. To illustrate the channel estimation process, we consider the $k$-th UE as an example. By accumulating spatial-channel samples over the medium-timescale, the average channel covariance matrix at the $t_{\mathrm{L}}$-th super-frame is given by
$\mathbf{R}_{k}^{(t_{\mathrm{L}})}=\frac{1}{t_{\mathrm{L}}T_{\mathrm{M}}}\sum_{t_{\mathrm{l}}=1}^{t_{\mathrm{L}}}\sum_{t_{\mathrm{M}}=1}^{T_{\mathrm{M}}}\widetilde{\mathbf{h}}^{(t)}_{\mathrm{S},k}(\widetilde{\mathbf{h}}_{\mathrm{S},k}^{(t)})^{H},~{}\forall k,t_{\mathrm{L}},$ (17)
where $t=(t_{\mathrm{l}}-1)T_{\mathrm{M}}+t_{\mathrm{M}}$ is defined for brevity and $\widetilde{\mathbf{h}}_{\mathrm{S},k}$ represents the estimated channel excluding the influence of the radiation patterns, given by
$\widetilde{\mathbf{h}}_{\mathrm{S},k}=\sum_{l=1}^{L}\alpha_{k,l}\mathbf{a}_{k,l},~{}\forall k.$ (18)
Then, the APS can be derived as
$\rho^{(t_{\mathrm{L}})}_{k}(\theta)=\frac{1}{\mathbf{e}^{H}(\theta)(\mathbf{R}_{k}^{(t_{\mathrm{L}})})^{-1}\mathbf{e}(\theta)},~{}\forall k,t_{\mathrm{L}},$ (19)
with
$\mathbf{e}(\theta)\triangleq[1,e^{-\jmath\frac{2\pi}{\lambda}d\sin\theta},\ldots,e^{-\jmath\frac{2\pi}{\lambda}(N_{\mathrm{t}}-1)d\sin\theta}]^{T}\in\mathbb{C}^{N_{\mathrm{t}}}.$ (20)
Since the APS reflects the power distribution across the angular domain, each peak in the APS physically represents the presence of a distinct scattering cluster in the propagation environment. Thus, the number of the peaks in $\rho^{(t_{\mathrm{L}})}_{k}(\theta)$ that exceed a specified threshold directly provides the estimated number of scattering clusters $C$. For each of these peaks, the corresponding angle value and power value indicate the estimated nominal angle $\overline{\theta}_{k,c}$ and the average channel gain $\alpha_{k,c}$ of the corresponding scattering cluster, respectively. In conclusion, referring to (12), we can construct the EM domain channel, which serves as the foundation for the long-timescale EM domain radiation beamformer design, which yields
$\mathbf{h}_{\mathrm{C},k}=\overline{\mathbf{H}}_{\mathrm{EM},k}\overline{\mathbf{h}}_{\mathrm{S},k}\in\mathbb{C}^{MN_{\mathrm{t}}},$ (21)
where the term $\overline{\mathbf{h}}_{\mathrm{S},k}\in\mathbb{C}^{CN_{\mathrm{t}}}$ encapsulates the spatial domain channel characteristics of $C$ scattering cluster cores and $\mathbf{\overline{H}}_{\mathrm{EM},k}\in\{0,1\}^{MN_{\mathrm{t}}\times CN_{\mathrm{t}}}$ indicates the angle information of the scattering clusters.
III-B.2 Long-timescale Radiation Beamformer Design
Notably, over the long-timescale, the EM domain radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ mainly aims to establish a basic pattern alignment with the scattering cluster core. This optimization requires only the channel information of the scattering cluster core, i.e., the estimated nominal angles and average channel gains of the scattering clusters. Consequently, we substitute the instantaneous multi-path channel $\mathbf{h}_{k}$ in the SINR expression (14) and the sum-rate formulation (15) with the scattering cluster core channel $\mathbf{h}_{\mathrm{C},k}$, defining the long-timescale sum-rate $\overline{R}_{k}$ as our optimization objective. Given a fixed analog beamformer $\mathbf{F}_{\mathrm{RF}}$ and digital beamformer $\mathbf{F}_{\mathrm{BB}}$, the long-timescale EM domain radiation beamforming design problem is then formulated as
$\max_{\mathbf{F}_{\mathrm{EM}}}$ $\ \ \sum^{K}_{k=1}\overline{R}_{k}$ (22a)
$\mathrm{s.t.}$ $\ \ \|\mathbf{f}_{\mathrm{EM},n}\|^{2}=1,~{}\forall n.$ (22b)
Considering the hardware implementability of pixel-based pattern-reconfigurable antennas, we adopt a practical approach by selecting the optimal radiation pattern from a predefined set, rather than designing arbitrary patterns with unrestricted directions and shapes. This approach ensures a balance between implementation efficiency and real-world hardware constraints. To be specific, the EM domain radiation beamformer is formulated as
$\mathbf{F}_{\mathrm{EM}}\triangleq\overline{\mathbf{F}}_{\mathrm{pat}}\mathbf{S}_{\mathrm{EM}},$ (23)
where $\overline{\mathbf{F}}_{\mathrm{pat}}\triangleq\mathbf{I}_{N_{\mathrm{t}}}\otimes\mathbf{F}_{\mathrm{pat}}\in\mathbb{R}^{MN_{\mathrm{t}}\times PN_{\mathrm{t}}}$ with $\mathbf{F}_{\mathrm{pat}}\in\mathbb{R}^{M\times P}$ representing a predefined radiation pattern dictionary that contains the radiation gains of $P$ candidate patterns over $M$ spatial sampling angles. The matrix $\mathbf{S}_{\mathrm{EM}}\triangleq\mathrm{blkdiag}\{\mathbf{s}_{\mathrm{EM},1},\mathbf{s}_{\mathrm{EM},2},\ldots,\mathbf{s}_{\mathrm{EM},N_{\mathrm{t}}}\}\in\{0,1\}^{PN_{\mathrm{t}}\times N_{\mathrm{t}}}$ denotes the EM domain pattern selection matrix, where each $\mathbf{s}_{\mathrm{EM},n}\in\{0,1\}^{P}$ selects a single optimal pattern from the dictionary for the $n$-th antenna, i.e., $\|\mathbf{s}_{\mathrm{EM},n}\|_{1}=1$. Based on this formulation, the original problem in (22) is equivalently transformed into a more practical optimization problem focused on optimizing the EM-domain pattern selection matrix $\mathbf{S}_{\mathrm{EM}}$, given by
$\max_{\mathbf{S}_{\mathrm{EM}}}$ $\ \ \sum^{K}_{k=1}\overline{R}_{k}$ (24a)
$\mathrm{s.t.}$ $\ \ \mathbf{s}_{\mathrm{EM},n}(p)\in\{0,1\},~{}\forall n,p,$ (24b)
$\ \ \|\mathbf{s}_{\mathrm{EM},n}\|_{1}=1,~{}\forall n.$ (24c)
The non-convex optimization problem in (24) poses significant challenges due to the fractional and logarithmic terms in the objective function (24a), as well as the non-smooth, non-convex Boolean constraint in (24b). To address these difficulties, we first apply the fractional programming (FP) method to reformulate the objective function (24a) into a polynomial expression, making it more tractable. Subsequently, we transform the Boolean constraint (24b) into a quadratic penalty term with a box constraint and leverage the majorization-minimization (MM) method to facilitate efficient optimization.
Firstly, by employing the Lagrangian dual reformulation and introducing auxiliary variable $\bm{\mu}\triangleq[\mu_{1},\mu_{2},\ldots,\mu_{K}]^{T}$, the objective function (24) can be transformed to
$\sum_{k=1}^{K}\log_{2}(1+\mu_{k})-\sum_{k=1}^{K}\mu_{k}$ (25)
$\ \ \ \ \ +\sum_{k=1}^{K}\frac{(1+\mu_{k})|\mathbf{h}_{\mathrm{C},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},k}|^{2}}{\sum_{j=1}^{K}|\mathbf{h}_{\mathrm{C},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},j}|^{2}+\sigma_{k}^{2}},$
which is equivalent to the original objective function (24a) when the auxiliary variable $\bm{\mu}$ has the optimal value
$\mu_{k}=\frac{|\mathbf{h}_{\mathrm{C},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},k}|^{2}}{\sum^{K}_{j=1,j\neq k}|\mathbf{h}_{\mathrm{C},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},j}|^{2}+\sigma_{k}^{2}},~{}\forall k.$ (26)
However, the summation of fractional terms in (25) still hinders a straightforward solution. Thus, we further adopt the quadratic transform to convert it into
$2\sqrt{1+\mu_{k}}\Re\{\xi_{k}^{\ast}\mathbf{h}_{\mathrm{C},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},k}\}-|\xi_{k}|^{2}D_{k},~{}\forall k,$ (27)
where $\bm{\xi}\triangleq[\xi_{1},\xi_{2},\ldots,\xi_{K}]^{T}$ is the auxiliary variable and $D_{k}\triangleq\sum_{j=1}^{K}|\mathbf{h}_{\mathrm{C},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},j}|^{2}+\sigma_{k}^{2}$ is utilized to simplify the expression. The expression (27) is equivalent to the last fractional term in (25) when the auxiliary variable $\bm{\xi}$ has the optimal value
$\xi_{k}=\frac{\sqrt{1+\mu_{k}}\mathbf{h}_{\mathrm{C},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},k}}{D_{k}},~{}\forall k.$ (28)
Based on the above derivation, the objective function can be reformulated as
${\sum_{k=1}^{K}}\big{[}\log_{2}(1+\mu_{k})-\mu_{k}$ (29)
$\ +2\sqrt{1+\mu_{k}}\Re\{\xi_{k}^{\ast}\mathbf{h}_{\mathrm{C},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},k}\}-|\xi_{k}|^{2}D_{k}\big{]}.$
In order to facilitate the subsequent transformation, (29) can be rewritten as the following concise form
$\sum_{k=1}^{K}\big{[}\log_{2}(1+\mu_{k})-\mu_{k}-|\xi_{k}|^{2}\sigma_{k}^{2}\big{]}+\delta,$ (30)
where we define
$\delta\triangleq$ $\sum_{k=1}^{K}\big{[}2\sqrt{1+\mu_{k}}\Re\{\xi_{k}^{\ast}\mathbf{h}_{\mathrm{C},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},k}\}$ (31)
$\hskip 28.45274pt-|\xi_{k}|^{2}\sum_{j=1}^{K}|\mathbf{h}_{\mathrm{C},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},j}|^{2}\big{]}.$
Having the optimal auxiliary variables $\bm{\mu}$ and $\bm{\xi}$, maximizing (30) is equivalent to maximizing the term $\delta$. Consequently, we can recast the optimization problem as
$\max_{\mathbf{S}_{\mathrm{EM}}}$ $\ \ \ \delta$ (32a)
$\mathrm{s.t.}$ $\ \ \ \mathbf{s}_{\mathrm{EM},n}(p)\in\{0,1\},~{}\forall n,p,$ (32b)
$\ \ \ \|\mathbf{s}_{\mathrm{EM},n}\|_{1}=1,~{}\forall n.$ (32c)
After dealing with the objective function, we turn to tackle the neither smooth nor convex Boolean constraint (32b). As stated in [30], the Boolean constraint (32b) can be transformed into an optimization problem of maximizing a quadratic term with a box constraint, which yields
$\max_{\mathbf{S}_{\mathrm{EM}}}$ $\ \ \mathbf{s}_{\mathrm{EM}}^{T}(\mathbf{s}_{\mathrm{EM}}-\mathbf{1})$ (33a)
$\mathrm{s.t.}$ $\ \ \mathbf{s}_{\mathrm{EM},n}(p)\in[0,1],~{}\forall n,p,$ (33b)
where $\mathbf{s}_{\mathrm{EM}}\triangleq[\mathbf{s}_{\mathrm{EM},1}^{T},\mathbf{s}_{\mathrm{EM},2}^{T},\ldots,\mathbf{s}_{\mathrm{EM},N_{\mathrm{t}}}^{T}]^{T}\in\mathbb{R}^{PN_{\mathrm{t}}}$ is introduced for brevity and the binary variable $\mathbf{S}_{\mathrm{EM}}$ has been relaxed to be continuous. Then, we can incorporate the objective function (33a) as a penalty term into the problem (32) with a box constraint (33b), which is given by
$\max_{\mathbf{S}_{\mathrm{EM}}}$ $\ \ \ \delta+\varrho_{1}\mathbf{s}_{\mathrm{EM}}^{T}(\mathbf{s}_{\mathrm{EM}}-\mathbf{1})$ (34a)
$\mathrm{s.t.}$ $\ \ \ \mathbf{s}_{\mathrm{EM},n}(p)\in[0,1],~{}\forall n,p,$ (34b)
$\ \ \ \|\mathbf{s}_{\mathrm{EM},n}\|_{1}=1,~{}\forall n,$ (34c)
where $\varrho_{1}$ is the penalty parameter. The newly constructed objective function (34a) contains a non-concave quadratic term related to the variable $\mathbf{S}_{\mathrm{EM}}$. Hence, we employ the MM approach in [31] to convert it into a linear form at the current point $\mathbf{s}_{\mathrm{EM}}^{(q)}$ in the $q$-th iteration, which can be expressed as
$\mathbf{s}_{\mathrm{EM}}^{T}(\mathbf{s}_{\mathrm{EM}}-\mathbf{1})\geq 2(\mathbf{s}_{\mathrm{EM}}^{(q)})^{T}\mathbf{s}_{\mathrm{EM}}-(\mathbf{s}_{\mathrm{EM}}^{(q)})^{T}\mathbf{s}_{\mathrm{EM}}^{(q)}-\mathbf{s}_{\mathrm{EM}}^{T}\mathbf{1}.$ (35)
Finally, the optimization problem for designing the EM domain pattern selection matrix is formulated as
$\max_{\mathbf{S}_{\mathrm{EM}}}$ $\ \ \ \delta+\varrho_{1}(2\mathbf{s}_{\mathrm{EM}}^{(q)}-\mathbf{1})^{T}\mathbf{s}_{\mathrm{EM}}$ (36a)
$\mathrm{s.t.}$ $\ \ \ \mathbf{s}_{\mathrm{EM},n}(p)\in[0,1],~{}\forall n,p,$ (36b)
$\ \ \ \|\mathbf{s}_{\mathrm{EM},n}\|_{1}=1,~{}\forall n,$ (36c)
which is convex and can be easily tackled by off-the-shelf solver CVX. With the optimal EM domain pattern selection matrix $\mathbf{S}_{\mathrm{EM}}$ and the known radiation pattern dictionary $\overline{\mathbf{F}}_{\mathrm{pat}}$, the radiation beamformer is constructed as $\mathbf{F}_{\mathrm{EM}}\triangleq\overline{\mathbf{F}}_{\mathrm{pat}}\mathbf{S}_{\mathrm{EM}}$ in (23).
III-C Medium-timescale RF Domain Design
III-C.1 Medium-timescale RF Domain Channel Estimation
The presence of reconfigurable antennas inherently couples the spatial-domain channel with the radiation pattern gains, as characterized in (12). Conventional channel estimation techniques designed for hybrid architectures do not account for this coupling, leading to estimation inaccuracies due to the neglect of radiation pattern effects on the measured channel [32], [33]. To address this limitation, it is essential to refine traditional estimation methods. Leveraging the inherent sparsity of the scattering cluster channel, we employ a compressed sensing (CS) algorithm for efficient channel estimation. Specifically, we construct a revised dictionary matrix that explicitly incorporates the impact of pattern-reconfigurable antennas, enabling a more accurate representation of the spatial-domain channel characteristics. This enhancement not only improves channel estimation accuracy but also facilitates a more robust and adaptive beamforming strategy, ultimately enhancing overall system performance.
Initially, the RF domain channel to be estimated is expressed as
$\mathbf{h}_{\mathrm{RF},k}\triangleq\mathbf{F}_{\mathrm{EM}}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}\overline{\mathbf{h}}_{\mathrm{S},k},~{}\forall k,$ (37)
where $\mathbf{F}_{\mathrm{EM}}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}$ reflects the radiation gains of the reconfigurable antennas based on the fixed radiation beamformer determined in the long-timescale, and it is apparent that this term is coupled with the spatial channel $\overline{\mathbf{h}}_{\mathrm{S},k}$. During the training phase, all UEs simultaneously transmit pilot signals on distinct frequency tones to the BS in the uplink. Taking the $k$-th UE as an example, the received signal at the BS is given by
$\mathbf{y}_{k}$ $=\sqrt{P_{\mathrm{t}}}\mathbf{W}_{k}^{H}\mathbf{h}_{\mathrm{RF},k}+\mathbf{W}_{k}^{H}\mathbf{n}_{k}$ (38)
$=\sqrt{P_{\mathrm{t}}}\mathbf{W}_{k}^{H}\mathbf{A}_{\mathrm{D}}\mathbf{z}_{k}+\mathbf{W}_{k}^{H}\mathbf{n}_{k}$
$=\mathbf{Q}_{k}\mathbf{z}_{k}+\mathbf{W}_{k}^{H}\mathbf{n}_{k},~{}\forall k,$
where $\mathbf{W}_{k}\in\mathbb{C}^{N_{\mathrm{t}}\times N_{\mathrm{Beam}}}$ is the training beamformer with $N_{\mathrm{Beam}}$ denoting the number of training beams, $\mathbf{A}_{\mathrm{D}}\in\mathbb{C}^{N_{\mathrm{t}}\times M}$ is the refined dictionary matrix incorporating the effects of pattern-reconfigurable antennas, $\mathbf{z}_{k}$ is a sparse vector with non-zero elements corresponding to the spatial angles, and $\mathbf{Q}_{k}\triangleq\sqrt{P_{\mathrm{t}}}\mathbf{W}_{k}^{H}\mathbf{A}_{\mathrm{D}}$ is the sensing matrix.
The conventional dictionary matrix consists of a set of steering vectors corresponding to uniformly distributed candidate angles, given by
$\mathbf{E}=[\mathbf{e}(\theta_{1}),\mathbf{e}(\theta_{2}),\ldots,\mathbf{e}(\theta_{M})]\in\mathbb{C}^{N_{\mathrm{t}}\times M},$ (39)
where each column $\mathbf{e}(\theta_{m})$ represents a steering vector pointing toward the angle $\theta_{m}=\pi m/M,~{}m=1,2,\ldots,M$. To explicitly account for the influence of the reconfigurable antenna pattern on the channel representation, we refine the dictionary matrix as
$\mathbf{A}_{\mathrm{D}}(:,m)=\mathbf{F}_{\mathrm{EM}}(:,m)\odot\mathbf{e}(\theta_{m}),~{}\forall m,$ (40)
where $\mathbf{F}_{\mathrm{EM}}(:,m)$ characterizes the radiation gains of the $N_{\mathrm{t}}$ antennas in the direction of $\theta_{m}$. This formulation effectively integrates both radiation pattern effects and angular information, enabling a more accurate estimation of spatial channel characteristics.
Subsequently, the channel estimation problem can be formulated as a sparse recovery problem:
$\min_{\mathbf{z}_{k}}$ $\ \ \ \|\mathbf{y}_{k}-\mathbf{Q}_{k}\mathbf{z}_{k}\|_{2}$ (41a)
$\mathrm{s.t.}$ $\ \ \ \|\mathbf{z}_{k}\|_{0}=L.$ (41b)
This problem is typically solved using the orthogonal matching pursuit (OMP) algorithm, which iteratively selects the most relevant dictionary elements to approximate the signal. The iterations continue until the residual error falls below a predefined threshold or the sparsity constraint is met. The implementation details of the OMP algorithm and the design of the training hybrid beamformer follow the principles outlined in [32], [33]. For brevity, these details are omitted here.
III-C.2 Medium-timescale Analog Beamformer Design
With the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ determined in the long-timescale, we now focus on designing the analog beamformer in the medium-timescale based on the statistical CSI of multi-path rays. Given the statistical information of the channel, the objective function is adjusted to maximize the ergodic sum-rate, ensuring robust performance over average channel conditions. Thus, the analog beamformer design problem in the medium-timescale is formulated as follows
$\max_{\mathbf{F}_{\mathrm{RF}}}$ $\ \ \sum^{K}_{k=1}\widetilde{R}_{k}$ (42a)
$\mathrm{s.t.}$ $\ \ \mathbf{F}_{\mathrm{RF}}(n,l_{\mathrm{RF}})\in\{\mathcal{F},0\},~{}\forall n,l_{\mathrm{RF}},$ (42b)
$\ \ \|\mathbf{F}_{\mathrm{RF}}(n,:)\|_{0}=1,~{}\forall n,$ (42c)
where (42a) defines the ergodic sum-rate $\widetilde{R}_{k}\triangleq\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{R_{k}\}$, evaluated over the RF domain channel samples, denoted as $\mathbf{H}_{\mathrm{RF}}\triangleq[\mathbf{h}_{\mathrm{RF},1},\mathbf{h}_{\mathrm{RF},2},\ldots,\mathbf{h}_{\mathrm{RF},K}]\in\mathbb{C}^{N_{\mathrm{t}}\times K}$ with $\mathbf{h}_{\mathrm{RF},k}$ given in (37). Notably, in this RF domain analog beamformer design problem, the power constraint is not explicitly imposed since the digital beamformer can be subsequently adjusted to satisfy it.
Unlike conventional statistical CSI-based beamforming approaches that require explicit knowledge of the channel statistics, we adopt an online learning framework that incrementally updates the channel statistical information by observing one channel sample at a time. This learning-based approach significantly reduces measurement overhead and latency, making it particularly well-suited for the tri-hybrid beamforming architecture. However, solving (42) is challenging due to the combination of discrete phase constraints in (42b) and the sparsity constraint in (42c), which collectively impose a highly non-convex set for $\mathbf{F}_{\mathrm{RF}}$. To overcome these challenges, we decompose the analog beamformer into two components:
$\mathbf{F}_{\mathrm{RF}}\triangleq\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}},$ (43)
where $\mathbf{S}_{\mathrm{RF}}\in\{0,1\}^{N_{\mathrm{t}}\times N_{\mathrm{RF}}2^{B}}$ is the RF domain phase-shift selection matrix and $\mathbf{F}_{\mathrm{set}}$ is defined as $\mathbf{F}_{\mathrm{set}}\triangleq\mathbf{I}_{N_{\mathrm{RF}}}\otimes\mathbf{f}_{\mathrm{set}}$ with
$\mathbf{f}_{\mathrm{set}}\triangleq\frac{1}{\sqrt{N_{\mathrm{t}}}}\left[1,e^{-\jmath\frac{2\pi}{2^{B}}},\ldots,e^{-\jmath\frac{2\pi(2^{B}-1)}{2^{B}}}\right]^{T},$ (44)
containing all possible phase-shifts for a low-resolution PS. The binary matrix $\mathbf{S}_{\mathrm{RF}}$ encodes phase-shift selections, where $\mathbf{S}_{\mathrm{RF}}(n,l_{\mathrm{RFB}})=1$ indicates that the $n$-th antenna is assigned to the $l_{\mathrm{RFB}}/2^{B}$-th RF chain with the corresponding phase-shift value given by $\mathbf{F}_{\mathrm{set}}(l_{\mathrm{RFB}},l_{\mathrm{RFB}}/2^{B})$. With $\mathbf{F}_{\mathrm{set}}$ predefined, the analog beamformer design problem (42) can be equivalently formulated as
$\max_{\mathbf{S}_{\mathrm{RF}}}$ $\ \ \ \sum^{K}_{k=1}\widetilde{R}_{k}$ (45a)
$\mathrm{s.t.}$ $\ \ \ \mathbf{S}_{\mathrm{RF}}(n,l_{\mathrm{RFB}})\in\{0,1\},~{}\forall n,l_{\mathrm{RFB}},$ (45b)
$\ \ \ \|\mathbf{S}_{\mathrm{RF}}(n,:)\|_{1}=1,~{}\forall n,$ (45c)
where the sum-rate expression in (45a) can be recast as
$\widetilde{R}_{k}=\log(1+\frac{|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},k}|^{2}}{\sum_{j=1,j\neq k}^{K}|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},j}|^{2}+\sigma_{k}^{2}}),~{}\forall k.\$ (46)
The primary challenges in solving this problem stem from the stochastic nature of the objective function (45a), which lacks a closed-form expression, and the non-smooth, non-convex Boolean constraint (45b). To address these issues, we employ a stochastic successive convex approximation (SSCA) framework [25], which iteratively refines the beamformer based on observed channel realizations. The expectation in the objective function is approximated using a quadratic surrogate function, allowing for online adaptation without requiring explicit statistical knowledge of the channel. Furthermore, the Boolean constraint is handled using a penalty-based approach, similar to the method described in Section III-B.
First, we construct a recursive quadratic surrogate function for the objective function (45a), leveraging both current and past channel samples. This iterative approach enables the BS to gradually learn the statistical properties of the channel by continuously accumulating channel observations, thereby enhancing beamforming performance over time and ensuring convergence. For convenience in the following expressions, we define $g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\triangleq R_{k}$, and introduce $\mathbf{s}_{\mathrm{RF}}\triangleq\mathrm{vec}(\mathbf{S}_{\mathrm{RF}})$.
At the $t_{\mathrm{M}}$-th recursion, based on the acquired channel sample $\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$ and the previous analog beamformer $\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$, we can substitute (45a) with a quadratic surrogate function as follows
$f^{(t_{\mathrm{M}})}(\mathbf{s}_{\mathrm{RF}})\!=\!v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}+(\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})})^{T}(\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})})-\tau\|\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}\|^{2},$ (47)
where $\tau>0$ is a constant ensuring strong convexity of $\tau\|\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}\|^{2}$. The term $v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}$ serves as an approximation of the objective function value $\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\}$, which is updated recursively as
$v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}=(1-\eta^{(t_{\mathrm{M}})})v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}}-1)}+\eta^{(t_{\mathrm{M}})}g_{0}(\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})};\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}),$ (48)
where $\eta^{(t_{\mathrm{M}})}\in(0,1]$ is a properly chosen sequence to satisfy convergence requirements. Similarly, the gradient term $\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}$ is an approximation of the partial derivative of $\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\}$ with respect to $\mathbf{s}_{\mathrm{RF}}$, evaluated at $\mathbf{s}_{\mathrm{RF}}=\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$. This is recursively computed as
$\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}=(1-\eta^{(t_{\mathrm{M}})})\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}}-1)}+\eta^{(t_{\mathrm{M}})}\nabla_{\mathbf{s}_{\mathrm{RF}}}g_{0}(\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})};\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}).$ (49)
The gradient of $g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})$ with respect to $\mathbf{s}_{\mathrm{RF}}$ is given by
$\nabla_{\mathbf{s}_{\mathrm{RF}}}g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})=\sum_{k=1}^{K}\left(\frac{\sum_{i}^{K}\mathbf{e}_{k,i}}{\mit{\Gamma_{k}}}-\frac{\sum_{i\neq k}^{K}\mathbf{e}_{k,i}}{\mit{\Gamma_{-k}}}\right),$ (50)
where we define
$\mit{\Gamma_{k}}\!$ $=\!\sum_{i}^{K}\left|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},i}\right|^{2}+\sigma_{k}^{2},~{}\forall k,$ (51a)
$\mit{\Gamma_{-k}}\!$ $=\!\sum_{i\neq k}^{K}\left|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},i}\right|^{2}+\sigma_{k}^{2},~{}\forall k,$ (51b)
$\mathbf{e}_{k,i}\!$ $=2\Re\{\mathrm{vec}((\mathbf{h}_{\mathrm{RF},k}\mathbf{h}_{\mathrm{RF},k}^{H})\mathbf{S}_{\mathrm{RF}}(\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},i}\mathbf{f}_{\mathrm{BB},i}^{H}\mathbf{F}_{\mathrm{set}}^{H}))\}.$ (51c)
Due to space limitations, we omit the detailed derivations here. The surrogate function $f^{(t_{\mathrm{M}})}(\mathbf{s}_{\mathrm{RF}})$ in (47) provides a concave approximation of the original stochastic objective function $\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\}$ in (45a), enabling efficient processing of problem (45) without explicitly computing the expectation.
To deal with the Boolean constraint (45b), we employ the same technique as specified in (33)-(35). As a result, the analog beamformer design problem is formulated as
$\max_{\mathbf{S}_{\mathrm{RF}}}$ $\ \ \ f^{(t_{\mathrm{M}})}(\mathbf{s}_{\mathrm{RF}})+\varrho_{2}(2\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}-\mathbf{1})^{T}\mathbf{s}_{\mathrm{RF}}$ (52a)
$\mathrm{s.t.}$ $\ \ \ \mathbf{S}_{\mathrm{RF}}(n,l_{\mathrm{RFB}})\in[0,1],~{}\forall n,l_{\mathrm{RFB}},$ (52b)
$\ \ \ \|\mathbf{S}_{\mathrm{RF}}(n,:)\|_{1}=1,~{}\forall n,$ (52c)
in which $\varrho_{2}$ denotes the penalty parameter. The resulting problem can be efficiently solved using a variety of existing methods or convex optimization solvers. Once the optimal solution, denoted as $\mathbf{\overline{S}}^{(t_{\mathrm{M}})}_{\mathrm{RF}}$, is obtained, the phase-shift selection matrix $\mathbf{S}_{\mathrm{RF}}$ is updated to incorporate both the newly computed solution and the previous information. This update is performed using the following recursive expression:
$\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}=(1-\eta^{(t_{\mathrm{M}})})\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}+\eta^{(t_{\mathrm{M}})}\mathbf{\overline{S}}^{(t_{\mathrm{M}})}_{\mathrm{RF}}.$ (53)
With the updated phase-shift selection matrix $\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}$, the analog beamformer is then constructed as $\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}=\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}\mathbf{F}_{\mathrm{set}}$.
III-D Short-timescale Digital Beamformer Design
In the $t_{\mathrm{S}}$-th short time-slot of the $t_{\mathrm{M}}$-th frame, the digital beamformer is updated while keeping the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ and the analog beamformer $\mathbf{F}_{\mathrm{RF}}$ fixed. The update relies on the instantaneous low-dimensional effective CSI, denoted as $\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})}=(\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{S}})})^{H}\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}})}\in\mathbb{C}^{K\times N_{\mathrm{RF}}}$, which is obtained from the processing of pilot signals in the BB domain. To design the digital beamformer, we apply the classic minimum mean square error (MMSE) method. The resulting digital beamformer is given by
$\mathbf{\widetilde{F}}_{\mathrm{BB}}^{(t_{\mathrm{S}})}=(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})})^{H}\left(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})}(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})})^{H}+\sigma_{k}^{2}\mathbf{I}_{K}\right)^{-1}.$ (54)
Finally, to ensure compliance with the total power constraint, the digital beamformer is normalized as
$\mathbf{F}_{\mathrm{BB}}^{(t_{\mathrm{S}})}=\frac{\sqrt{P_{\mathrm{t}}}\ \mathbf{\widetilde{F}}_{\mathrm{BB}}^{(t_{\mathrm{S}})}}{\|\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}})}\mathbf{\widetilde{F}}_{\mathrm{BB}}^{(t_{\mathrm{S}})}\|_{F}}.$ (55)
Algorithm 1 The proposed tri-timescale tri-hybrid beamforming design algorithm.
  Initialize $\mathbf{F}_{\mathrm{EM}},\mathbf{F}_{\mathrm{RF}},\mathbf{F}_{\mathrm{BB}}$.
  Step 1: Long-timescale EM domain design at the $t_{\mathrm{L}}$-th super-frame
  • High-dimensional scattering cluster core channel estimation
    • Calculate the channel covariance matrix $\mathbf{R}_{k}^{(t_{\mathrm{L}})},\forall k$ by (17).
    • Derive the APS $\rho^{(t_{\mathrm{L}})}_{k}(\theta),\forall k$ as (19).
    • Estimate the nominal angles $\theta_{k,c},\forall k,c,$ and average channel gains $\alpha_{k,c},\forall k,c,$ of the scattering clusters.
    • Construct the EM domain channel $\mathbf{h}^{(t_{\mathrm{L}})}_{\mathrm{C},k}$ by (21).
  • Radiation beamformer design
    • Calculate EM domain pattern selection matrix $\mathbf{S}_{\mathrm{EM}}$.
      • while no convergence of $\mathbf{S}_{\mathrm{EM}}$ do
      • Update $\bm{\mu}$ by (26).
      • Update $\bm{\xi}$ by (28).
      • Update pattern selection matrix $\mathbf{S}_{\mathrm{EM}}$ by solving (36).
      • end while
    • Construct the radiation beamformer $\mathbf{F}^{(t_{\mathrm{L}})}_{\mathrm{EM}}$ by (23).
  Step 2: Medium-timescale RF domain design at the $t_{\mathrm{M}}$-th frame
  • Statistical high-dimensional multi-path rays channel estimation
    • Construct the modified dictionary matrix $\mathbf{A}^{(t_{\mathrm{M}})}_{\mathrm{D}}$ by (40).
    • Solve the channel estimation problem (41) by the OMP algorithm.
    • Construct the RF domain channel $\mathbf{h}^{(t_{\mathrm{M}})}_{\mathrm{RF},k},\forall k,$ by (37).
  • Analog beamformer design
    • Update the surrogate function (47) based on $\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})},\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$.
    • Calculate the optimal $\mathbf{\overline{S}}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$ using (52).
    • Update $\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}$ according to (53).
    • Construct the analog beamformer $\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}$ by (43).
  Step 3: Short-timescale BB domain design at the $t_{\mathrm{S}}$-th time-slot
  • Instantaneous low-dimensional effective channel estimation
    • Observe an instantaneous low-dimensional effective channel sample $\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})}$.
  • Digital beamformer design
    • With fixed $\mathbf{F}_{\mathrm{EM}}^{(t_{\mathrm{L}})}$ and $\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$, calculate the digital beamformer $\mathbf{F}_{\mathrm{BB}}^{(t_{\mathrm{S}})}$ according to (54) and (55).
III-E Summary, Complexity and Overhead Analysis
III-E.1 Summary
Based on the above derivations, the proposed tri-timescale tri-hybrid beamforming design is straightforward and summarized in Algorithm 1. After appropriately initializing the beamforming matrices $\mathbf{F}_{\mathrm{EM}}$, $\mathbf{F}_{\mathrm{RF}}$ and $\mathbf{F}_{\mathrm{BB}}$ across the three domains, the algorithm iteratively acquires the corresponding channel information and updates these matrices over the three timescales, respectively.
III-E.2 Complexity Analysis
Next, we provide a brief analysis of the algorithmic complexity. The primary computational complexity lies in calculating the EM domain pattern selection matrix $\mathbf{S}_{\mathrm{EM}}$ in (36) over the long-timescale, and determining the phase-shift selection matrix $\mathbf{S}_{\mathrm{RF}}$ in (52) over the medium-timescale. Specifically, the computational complexity of solving problem (36) in each super-frame over the long-timescale, which involves an $N_{\mathrm{t}}P$-dimensional variable and requires handling $2N_{\mathrm{t}}P$ second-order cone (SOC) constraints, is of order $\mathcal{O}\{\sqrt{2N_{\mathrm{t}}P+N_{\mathrm{t}}+1}N_{\mathrm{t}}P(N_{\mathrm{t}}^{2}P^{2}+2N_{\mathrm{t}}P+N_{\mathrm{t}})\}$ [34]. Similarly, the computational complexity of solving problem (52) in each frame during the medium-timescale, which entails calculating an $N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}$-dimensional variable along with $2N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}$ SOC constraints, is of order $\mathcal{O}\{\sqrt{2N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}+N_{\mathrm{t}}+1}N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}(N_{\mathrm{t}}^{2}N_{\mathrm{RF}}^{2}2^{2B}+2N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}+N_{\mathrm{t}})\}$.
III-E.3 Pilot Overhead Analysis
The pilot overhead in the proposed scheme primarily stems from channel sampling at the medium-timescale and short-timescale stages. In the medium-timescale, CS-based channel estimation requires $N_{\mathrm{Beam}}K/N_{\mathrm{RF}}$ pilot symbols to obtain the RF domain high-dimensional statistical channel sample $\mathbf{h}_{\mathrm{RF},k}^{(t_{\mathrm{M}})},\forall k$, as described in Section III-C.In the short-timescale, $N_{\mathrm{RF}}K$ pilot symbols are needed for estimating the BB domain low-dimensional effective channel $\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})}$, as described in Section III-D. Therefore, the total pilot overhead across the channel statistics coherence time is expressed as $(N_{\mathrm{Beam}}K/N_{\mathrm{RF}}+N_{\mathrm{RF}}KT_{\mathrm{S}})T_{\mathrm{M}}T_{\mathrm{L}}$. This is notably less than that of the real-timescale beamforming approaches, which requires $(N_{\mathrm{Beam}}K/N_{\mathrm{RF}}+N_{\mathrm{RF}}K)T_{\mathrm{S}}T_{\mathrm{M}}T_{\mathrm{L}}$.
IV Simulation Results
In this section, we provide simulation results to demonstrate the effectiveness of the proposed tri-timescale tri-hybrid beamforming design algorithm with reconfigurable antennas in enhancing communication performance while ensuring hardware efficiency. Unless otherwise specified, the BS is equipped with $N_{\mathrm{t}}=32$ antennas and the center carrier frequency is set to $f_{\mathrm{c}}=28$ GHz. The low-resolution PS used in the RF domain has quantized phase controlled by $B=3$ bits. Without loss of generality, we assume that the number of UEs $K$ is equal to the number of RF chains $N_{\mathrm{RF}}$. In the EM domain, each pixel-based reconfigurable antenna enables the generation of $P=7$ distinct directional radiation patterns by adjusting the electronic switches between pixels [8], with $M=180$ uniform sampling points in the azimuth dimension. In this scenario, there are $K=2$ UEs, all operating under a scattering cluster channel with $C=2$ clusters, each of which consists of $L_{c}=3$ multi-path rays. The nominal AoDs $\overline{\theta}_{k,c}$ of these scattering clusters are randomly distributed within $[-\pi/3,\pi/3]$, while the angle spread is set to $\varsigma_{k,c}=\pi/36$. The path loss is expressed as $\alpha_{k,c,l_{c}}=10^{-\frac{C_{0}}{10}}(\frac{r_{k,c,l_{c}}}{D_{0}})^{-\kappa_{k,c,l_{c}}}$ with $\kappa_{k,c,l_{c}}\in[2.5,3]$ being the attenuation factor, $C_{0}=30$ dB, $D_{0}=1$m. The propagation distance $r_{k,c,l_{c}}$ is composed of the nominal distance $\overline{r}_{k,c}\in[50\mathrm{m},100\mathrm{m}]$ and $\iota=0.5\mathrm{m}$. The noise power of UEs is set as $\sigma_{k}^{2}=-70$dBm. The tri-timescale framework is structured with $T_{\mathrm{L}}=10$ super-frames for the long-timescale, $T_{\mathrm{M}}=50$ frames for the medium-timescale, and $T_{\mathrm{S}}=200$ time-slots for the short-timescale.
(a) Long-timescale.
(b) Medium-timescale.
Figure 5: Convergence performance of the long-timescale radiation beamformer design and medium-timescale analog beamformer design.
The convergence performance of the proposed tri-timescale tri-hybrid beamforming design is illustrated in Fig. 5, which includes both the long-timescale radiation beamformer design algorithm and the medium-timescale analog beamformer design algorithm. It should be emphasized that the convergence implications differ between the two algorithms. In the long-timescale case, convergence refers to the iterative calculation of the radiation beamformer at the start of super-frame, based on the estimated CSI of the scattering cluster core. In contrast, for the medium-timescale case, the convergence process refers to the recursive learning of the channel statistics of the multi-path rays within the scattering clusters using high-dimensional channel samples and the recursive updating of the analog beamformer in each frame. The results presented in Fig. 5 demonstrate that both algorithms exhibit satisfactory convergence speed, indicating an efficient level of computational complexity.
Figure 6: The sum-rate versus the transmit power $P_{\mathrm{t}}$.
Fig. 6 illustrates the sum-rate performance versus transmit power for systems employing reconfigurable and conventional antennas. Specifically, the proposed tri-hybrid beamforming architecture with reconfigurable antennas and a dynamic-subarray scheme is denoted as “RA, DS”. For comparison, hybrid beamforming using conventional antennas with a dynamic-subarray is labeled “CA, DS”, and the fixed-subarray architectures with reconfigurable and conventional antennas are denoted as “RA, FS” and “CA, FS”, respectively. Performance is evaluated for systems with $N_{\mathrm{t}}=32$ and $N_{\mathrm{t}}=16$ antennas. As shown in Fig. 6, the proposed “RA, DS” algorithm consistently outperforms its competitors. In particular, reconfigurable antennas yield a power gain of approximately 7 dB compared to conventional antennas, while the dynamic-subarray scheme provides an additional gain of about 2 dB over the fixed-subarray approach. These results confirm that the tri-hybrid beamforming architecture achieves superior communication performance by virtue of its enhanced flexibility in radiation pattern selection in the EM domain and subarray selection in the RF domain.
Figure 7: The sum-rate and channel estimation overhead versus the number of frames $T_{\mathrm{M}}$.
For the proposed tri-timescale tri-hybrid beamforming design, increasing the channel sampling rate in the medium-timescale can enhance sum-rate performance by enabling more frequent updates of channel statistics for multi-path rays and scattering cluster cores. However, this improvement comes at the cost of higher pilot overhead. To quantitatively evaluate the impact of the channel sampling rate on both system performance and pilot overhead, we gradually increase the number of frames per super-frame in the medium-timescale, effectively increasing the channel sampling rate, as illustrated in Fig. 7. By analyzing the pilot overhead expression derived in Section III-E and observing the trend in Fig. 7, we conclude that the proposed tri-timescale scheme significantly reduces pilot overhead compared to the real-time scheme, which requires high-dimensional CSI estimation in every time-slot. More importantly, as the number of frames increases, the sum-rate performance of the tri-timescale beamforming scheme steadily improves and eventually saturates at a satisfactory level, albeit with a monotonic increase in channel estimation overhead. In particular, setting the number of frames to $T_{\mathrm{M}}=50$ achieves a favorable balance between performance and overhead, a configuration adopted in the simulations.
Figure 8: The impact of channel estimation configurations on sum-rate performance.
The channel estimation algorithms in both the EM and RF domains are critical components of the proposed tri-timescale framework. Fig. 8 illustrates the impact of different channel estimation configurations on sum-rate performance under various system settings. The results indicate that, regardless of the number of propagation paths $L_{\mathrm{c}}$, the sum-rate achieved with estimated CSI is highly comparable to that with perfect CSI. This finding underscores the accuracy and robustness of our channel estimation algorithms within the proposed framework, even under complex propagation conditions. Moreover, even with a reduced angular resolution $N_{\mathrm{Beam}}$, the channel characteristics are reliably captured, and acceptable sum-rate performance is achieved, further validating the effectiveness and reliability of the proposed approach.
Figure 9: The sum-rate versus the distance and angle of UE.
Finally, we validate the effectiveness of the proposed scheme by evaluating the impact of user location on the overall performance of the tri-hybrid beamforming with reconfigurable antennas (RA) and the hybrid beamforming with conventional antennas (CA). Specifically, Fig. 9 (a) depicts the sum-rate performance as a function of the nominal distance, which represents the distance between the BS and the UEs, while Fig. 9 (b) illustrates the sum-rate variation with respect to the nominal angle, which represents the user’s angular position relative to the antenna array. From Fig. 9 (a), we observe that as the nominal distance increases, the performance gain offered by pattern-reconfigurable antennas gradually diminishes. This is primarily due to the dominant effect of path loss, which attenuates signal power, reducing the relative impact of radiation pattern variations across different angles. In contrast, Fig. 9 (b) reveals that the RA-based scheme maintains robust performance across a wide range of user angles by dynamically adjusting the pattern orientations as the user’s position deviates from the optimal direction. In comparison, the CA-based scheme achieves satisfactory performance only when the user is positioned perpendicular to the array, where the antenna gain is maximized. These results underscore the robustness and versatility of the proposed tri-hybrid beamforming approach, demonstrating that dynamic pattern reconfiguration not only compensates for variations in user location but also substantially enhances overall system performance and reliability across a wide range of deployment scenarios.
V Conclusions
This paper presents a novel tri-hybrid beamforming architecture that integrates pattern-reconfigurable antennas with analog and digital beamforming. By leveraging an innovative tri-timescale beamforming framework, the proposed approach optimizes the radiation beamformer over a long-timescale, the analog beamformer over a medium-timescale, and the digital beamformer over a short-timescale. This hierarchical design effectively balances system performance with practical constraints, such as channel estimation overhead and computational complexity. Extensive simulation results validate the superiority of the proposed scheme over conventional hybrid beamforming architectures, highlighting its potential in future wireless communication systems.

### v2

@@
 This hierarchical strategy effectively balances performance and implementation feasibility.
-Simulation results validate the performance gains of the proposed tri-hybrid architecture and demonstrate that the tri-timescale design significantly reduces pilot overhead and computational complexity, highlighting its potential for future wireless communication systems.
-Index Terms:
-Reconfigurable antennas, tri-hybrid beamforming, tri-timescale, hybrid beamforming.
+Simulation results validate the performance gains of the proposed tri-hybrid architecture and demonstrate that the tri-timescale design significantly reduces pilot overhead and computational complexity, highlighting its potential for future wireless communication systems, and revealing a practical way to exploit EM‑domain reconfigurability across three natural channel timescales.
 I Introduction
-Future wireless communication systems are expected to leverage ultra-wide spectrum resources in high-frequency bands, such as millimeter wave (mmWave), to accommodate the growing demand for ultra-high data rates [1].
-To mitigate the severe propagation loss in these frequency bands, massive multiple-input multiple-output (mMIMO) systems or extremely large antenna arrays (ELAA) are typically employed to ensure reliable signal transmission [2].
-In conventional antenna arrays, each antenna element typically exhibits a fixed and identical radiation patterns that concentrate energy towards predetermined directions, often perpendicular to the array.
-Within this framework, most existing studies focus on signal processing-based beamforming techniques, which adjust the phase and amplitude of signals across antenna elements to direct beam towards specific directions.
-However, the limitations of conventional pattern-fixed antennas are becoming increasingly apparent in emerging applications, such as ELAA and near-field communications [3].
-Specifically, system performance is significantly influenced not only by array beamforming but also by the radiation gain of each individual antenna element.
-While antennas with narrow, highly energy-concentrated radiation patterns provide high gain within specific angular ranges, they suffer substantial performance degradation in the endfire direction due to a sharp decrease in signal strength.
-Conversely, antennas with broader radiation patterns provide wider coverage at the cost of reduced gain, ultimately compromising the quality of service.
-Reconfigurable antenna technology offers a promising solution to the challenges faced by conventional pattern-fixed antennas by enhancing adaptability [4]-[7].
-Reconfigurable antennas can adapt their internal parameters to modify key operating characteristics such as radiation patterns [8]-[10], frequency responses [11], and polarizations [12], [13].
-Among these, radiation pattern reconfiguration is particularly crucial for wireless communication systems, as it directly impacts both the transmitted and received signal strength.
-To enable flexible control over these characteristics, several hardware implementations have been developed, such as electronically steerable parasitic array radiator (ESPAR) antennas [14], frequency selective surfaces (FSS) [15], and pixel antennas [16], all of which share similar underlying principles and architectures.
-Typically, a reconfigurable antenna consists of an active element and multiple parasitic elements, where the connection states of parasitic elements are dynamically controlled to alter the current distribution within the antenna, thereby modifying its electromagnetic (EM) properties.
-This capability introduces new degrees of freedom (DoF) in the EM domain, offering the potential to overcome existing performance limitations and enable more efficient wireless communication systems.
-With significant advancements in reconfigurable antenna technology, integrating pattern-reconfigurable antennas into antenna arrays has gained considerable attention in recent years.
+Future wireless communication systems are anticipated to leverage ultra-wide spectrum resources in high-frequency bands, such as millimeter-wave (mmWave), to meet the escalating demand for ultra-high data rates [1].
+To counteract severe path loss in these bands, massive multiple-input multiple-output (massive MIMO) systems or extremely large antenna arrays (ELAA) have emerged as promising technologies to ensure reliable signal transmission [2].
+In conventional antenna arrays, each antenna element typically possesses a fixed radiation pattern, concentrating energy toward predetermined directions, often perpendicular to the array plane.
+Consequently, existing research primarily focuses on signal-processing-based beamforming methods that manipulate the phase and amplitude of antenna signals across antenna elements to steer beams toward specific directions.
+However, the limitations of conventional antennas with fixed radiation patterns become increasingly evident in emerging applications, such as ELAA and near-field communications [3].
+Specifically, system performance critically depends not only on array-level beamforming but also on individual antenna element radiation characteristics.
+Narrow radiation patterns with high energy concentration can deliver strong gains within specific angular sectors but experience significant performance deterioration in end-fire directions due to sharp gain reductions.
+Conversely, antennas with broader radiation patterns provide extended coverage at the expense of lower gain, ultimately degrading user quality of service (QoS).
+Reconfigurable antenna technology offers a compelling approach to addressing these limitations by introducing enhanced adaptability [4]-[9].
+Such antennas dynamically adjust their internal configurations, modifying key operational characteristics, including radiation patterns [10]-[12], frequency responses [13], and polarizations [14], [15].
+Among these, radiation pattern reconfiguration is particularly significant for wireless communication, directly influencing signal strength at transmission and reception.
+Several hardware implementations facilitating flexible control have emerged, including electronically steerable parasitic array radiator (ESPAR) antennas [16], frequency selective surfaces (FSS) [17], and pixel antennas [18].
+Typically, a reconfigurable antenna comprises an active radiating element and multiple parasitic elements, whose connection states are dynamically adjusted to alter current distribution and thereby modify electromagnetic (EM) properties.
+This adaptability provides new degrees of freedom (DoF) in the EM domain, offering substantial potential to overcome existing performance limitations and enable more efficient wireless communication systems.
+Recent advances in reconfigurable antenna technology have sparked growing interest in integrating pattern-reconfigurable antennas into arrays.
 For instance, the authors in [19] employ multi-port network theory and beamspace channel representation to develop an EM-based communication model for pixel antennas, which is then used to optimize antenna coding for maximizing channel gain.
-A similar approach is applied in [14] to construct an ESPAR antenna model, enabling analysis and optimization of ESPAR beamforming.
 Additionally, the authors in [20] introduce a decomposition method based on spherical harmonic functions to model radiation patterns in the EM domain.
 This approach decomposes the radiation pattern into a linear combination of orthogonal bases, facilitating the iterative design of radiation patterns and digital precoders.
-While these studies highlight the advantages of replacing conventional pattern-fixed antennas with innovative pattern-reconfigurable ones, they overlook the significant hardware costs and power consumption incurred by the increasing number of antennas and operating frequencies.
-To tackle these issues, existing literature commonly suggests hybrid beamforming architectures for mMIMO systems, as a power-efficient alternative to conventional fully-digital beamforming [19].
-In hybrid beamforming, analog beamforming controls the beam direction in the radio frequency (RF) domain, while digital beamforming performs spatial signal processing in the baseband (BB) domain, leveraging a reduced number of RF chains and hardware-efficient phase shifters (PSs).
-Common hybrid beamforming architectures, including fully-connected [20], partially-connected [21], and dynamic architectures [22], strike a balance between performance and implementation cost.
-While recent advances in hardware and power-efficient hybrid architectures have enabled large-scale MIMO deployments for 6G and beyond, the limitations of conventional pattern-fixed antennas remain a significant challenge.
-To address this issue, we propose leveraging the flexibility of pattern-reconfigurable antenna technology to address the performance degradation caused by fixed radiation patterns, while utilizing hybrid beamforming to reduce power consumption without sacrificing performance.
+While these studies highlight the advantages of replacing conventional pattern-fixed antennas with innovative pattern-reconfigurable ones, they often neglect the growing concerns of hardware complexity and energy consumption that future wireless communication systems will encounter.
+Such concerns become increasingly critical in large-scale deployments, where both the antenna count and operating frequency bands continue to expand.
+To tackle these issues, existing literature commonly advocates hybrid beamforming architectures for mMIMO systems, as a power-efficient alternative to conventional fully-digital beamforming [21].
+Hybrid beamforming architectures typically partition functionality between analog beamforming, controlling beam directions in the radio-frequency (RF) domain, and digital beamforming, performing spatial signal processing in the baseband (BB) domain.
+Utilizing fewer RF chains and hardware-efficient phase shifters (PSs), various hybrid architectures, such as fully-connected [22], partially-connected [23], and dynamic structures [24], achieve a balanced trade-off between performance and complexity.
+However, these architectures inherently rely on antennas with fixed radiation characteristics, severely restricting their adaptability to dynamic propagation environments and limiting overall system performance.
+To overcome this critical limitation, leveraging pattern-reconfigurable antennas within hybrid beamforming architectures emerges as an effective solution, enabling dynamic adaptation of radiation characteristics while maintaining reduced hardware complexity and power consumption.
 In this paper, we consider a novel tri-hybrid beamforming architecture that integrates pattern-reconfigurable antennas with both analog and digital beamforming.
-Specifically, this tri-hybrid architecture consists of three layers: A radiation beamformer in the EM domain to align radiation patterns with desired direction, an analog beamformer in the RF domain to enhance array gain, and a digital beamformer in the BB domain to mitigate multi-user interference.
-The concept of tri-hybrid architecture was first introduced in [23], [24].
-While prior analyses and simulations have demonstrated the advantages of tri-hybrid beamforming in terms of spectral efficiency and energy consumption, several critical issues must be addressed before its practical implementation in future wireless systems.
-First, accurate acquisition of instantaneous channel state information (CSI) becomes more challenging due to the higher dimensionality of CSI and the limited number of RF chains, leading to increased pilot overhead and requiring advanced estimation algorithms.
-Second, optimizing the beamformers across all three domains is more complicated due to parameter coupling, discrete phase-shift constraints, and the nonlinear characteristics of reconfigurable antennas, etc.
-Recently, a two-timescale framework has been proposed to address system overhead and design complexity in conventional hybrid beamforming.
-In this framework, the analog beamformer is optimized based on channel statistics over a long-timescale, while the digital beamformer is designed using low-dimensional instantaneous effective CSI over a short-timescale [25]-[27].
-Building on this concept and leveraging the unique characteristics of the tri-hybrid architecture, we propose a novel tri-timescale framework for tri-hybrid beamforming design, effectively managing practical overhead and complexity.
+This architecture can be viewed as an advanced extension of existing reconfigurable antenna-based frameworks.
+Specifically, the proposed tri-hybrid architecture comprises three distinct layers: an EM domain radiation beamformer aligning radiation patterns dynamically, an RF domain analog beamformer enhancing array gains, and a BB domain digital beamformer mitigating multi-user interference.
+The concept of this tri-hybrid architecture was first introduced in [25]-[27].
+While prior analyses and simulations have demonstrated the advantages of tri-hybrid beamforming in terms of spectral efficiency and energy efficiency, several critical issues must be addressed before its practical implementation in future wireless systems.
+Firstly, acquiring instantaneous channel state information (CSI) becomes considerably more challenging due to increased dimensionality and limited RF chains, resulting in substantial pilot overhead.
+Secondly, jointly optimizing beamformers across three distinct domains introduces considerable complexity, primarily due to parameter coupling, discrete phase-shift constraints, and the nonlinear characteristics inherent to reconfigurable antennas.
+To address these challenges, recent studies proposed a two-timescale beamforming framework to mitigate pilot overhead and computational complexity.
+Specifically, analog beamforming optimization leverages statistical CSI over a longer timescale, whereas digital beamforming employs instantaneous, low-dimensional effective CSI over a shorter timescale [28]-[30].
+Inspired by this approach and capitalizing on the distinctive properties of our tri-hybrid architecture, we introduce an innovative tri-timescale beamforming framework, effectively resolving practical implementation challenges, significantly reducing channel estimation overhead, and lowering computational complexity.
 The main contributions of this paper are summarized as follows:
-  • We first develop a comprehensive mathematical model for the radiation beamformer based on multiple pattern-reconfigurable antennas, along with the corresponding virtual angular index channel in the EM domain.
-Building upon this model, we derive the formulation of the tri-hybrid beamforming architecture, establish the signal model for a multi-user multiple-input single-output (MU-MISO) system employing this architecture, and present the associated optimization problem.
-  • For the tri-hybrid beamforming design, our objective is to maximize the sum-rate while satisfying constraints on radiation pattern energy, constant modulus and discrete phase shifts of PSs, a non-overlapping dynamic connection network, and the total transmit power budget.
-To effectively address the challenges of high pilot overhead and substantial computational complexity, we propose a novel tri-timescale beamforming framework, where the radiation beamformer is optimized over a long-timescale, the analog beamformer over a medium-timescale, and the digital beamformer over a short-timescale.
-  • Extensive simulations are conducted to validate the capability of the tri-hybrid beamforming architecture in enhancing system performance, as well as the effectiveness of the proposed tri-timescale beamforming design algorithm in reducing pilot overhead and computational complexity.
+  • We propose a comprehensive mathematical model for the radiation beamformer based on multiple pattern-reconfigurable antennas, including the introduction of a novel virtual angular index channel in the EM domain.
+Utilizing this model, we develop a tri-hybrid beamforming architecture, formulate a detailed multi-user multiple-input single-output (MU-MISO) signal model, and rigorously establish the corresponding sum-rate maximization optimization problem under realistic system constraints.
+  • To overcome the inherent challenges of excessive pilot overhead and high computational complexity, we introduce an innovative tri-timescale beamforming framework.
+Specifically, we optimize the radiation beamformer at a long timescale according to the statistical CSI of scattering cluster cores, update the analog beamformer at a medium timescale based on high-dimensional statistical CSI of multipath rays, and adaptively design the digital beamformer at a short timescale using instantaneous low-dimensional effective CSI.
+  • We conduct extensive simulations to verify the effectiveness and scalability of the proposed tri-hybrid beamforming architecture.
+Particularly, the proposed tri-hybrid beamforming achieves around 7dB transmit power reduction at a given sum-rate and nearly an order-of-magnitude pilot overhead reduction compared to real-time beamforming, underscoring its practical advantages for next-generation wireless systems.
 Figure 1: The hardware prototypes of pattern-reconfigurable antenna (left [11], right [12]).
@@
 II-A Pattern-reconfigurable Antenna
-Pattern-reconfigurable antenna is an advanced technology that enables dynamic adaptation of radiation patterns to enhance signal quality, mitigate interference, and improve coverage in wireless communication systems.
-Various hardware implementations have been explored to achieve flexible radiation pattern control, including ESPAR antennas [14], FSS [15], and pixel antennas [16].
-Among these, pixel antenna technology has emerged as the most practical and versatile solution for integration into compact antenna arrays.
-The hardware prototypes of pixel-based reconfigurable antennas are illustrated in Fig.
+Pattern-reconfigurable antennas represent advanced technologies that dynamically adapt their radiation patterns, significantly enhancing signal quality, mitigating interference, and improving coverage in wireless communication systems.
+Several hardware implementations have been explored to achieve such flexible radiation control, notably ESPAR antennas [16], FSS [17], and pixel antennas [18].
+Among these, pixel antenna technology has emerged as the most practical and versatile solution, particularly suitable for integration into compact antenna arrays.
+Hardware prototypes of pixel-based reconfigurable antennas are illustrated in Fig.
 1.
-Having demonstrated their feasibility in real-world applications, we now describe the hardware configuration and operational principle of the pixel-based pattern-reconfigurable antennas.
+This prototype employs PIN-diode switching networks that exhibit microsecond-level switching latencies and highly reliable reconfiguration behavior, confirming their suitability for rapid EM domain adaptation within the proposed framework.
+Moreover, these antennas typically require only a minimal number of serialized binary control lines and low-voltage direct-current (DC) bias circuits, resulting in affordable control signaling overhead and power consumption.
+Having established their practicality for real-world deployment, we now detail the hardware configuration and operating principles of pixel-based pattern-reconfigurable antennas.
+As depicted in Fig.
+2, pixel antenna technology discretizes a continuous antenna surface into numerous electrically small parasitic elements, termed pixels, driven centrally by an active radiating source.
+These pixels, functioning as fundamental building blocks, employ low-voltage, binary-controlled PIN diodes to connect or disconnect adjacent elements.
+By adjusting these switches’ states, various pixel topologies are formed, modifying the current distribution across the antenna surface.
+This reconfiguration process effectively alters the radiated electric and magnetic field distributions, thus enabling the generation of diverse radiation patterns.
+To rigorously evaluate the performance improvements brought by pattern-reconfigurable antennas, it is crucial to establish a mathematical model capturing both radiation patterns and corresponding channel characteristics.
+As aforementioned, pattern-reconfigurable antennas introduce an additional DoF in the EM domain, effectively expanding conventional antennas’ spatial representation from a one-dimensional model to an $M$-dimensional angular domain.
 As illustrated in Fig.
-2, pixel antenna technology discretizes a continuous antenna surface into electrically small parasitic elements, called pixels, with a central radiating source driving the entire antenna.
-These pixels serving as fundamental building blocks, utilize low-voltage, binary-controlled pin diodes to connect or disconnect adjacent pixels.
-By adjusting the states of these switches, different pixel topologies are formed, which modify the current distribution across the antenna surface.
-This reconfiguration alters the electric and magnetic field components radiated by the antenna, enabling the generation of a diverse set of radiation patterns.
-To evaluate the performance improvements brought by pattern-reconfigurable antennas, it is crucial to establish a mathematical model for both the radiation pattern and the corresponding channel characteristics.
-As aforementioned, pattern-reconfigurable antennas introduce an additional DoF in the EM domain, effectively expanding the information coverage dimension of conventional antennas, from a one-dimensional spatial representation to an $M$-dimensional angular information, where $M$ represents the number of uniform sampling points in the angular domain.
-Specifically, a single reconfigurable radiation pattern (exemplified by the red pattern in Fig.
-2) is characterized over $M$ discrete directions in the EM domain and is represented by $\mathbf{f}_{\mathrm{EM}}\in\mathbb{R}^{M}$, with each element corresponding to the radiation gain in a specific direction.
-To model reception characteristics, we introduce the virtual angular index channel for the pattern-reconfigurable antenna, denoted as $\mathbf{h}_{\mathrm{EM}}\in\{0,1\}^{M}$, which serves as an index matrix indicating the spatial angles at which the user is located.
-Consequently, the effective gain in the EM domain can be expressed as $G_{\mathrm{EM}}=\mathbf{h}_{\mathrm{EM}}^{H}\mathbf{f}_{\mathrm{EM}}$, which quantifies the radiation intensity directed toward the user’s actual spatial angle.
-By optimizing the radiation pattern $\mathbf{f}_{\mathrm{EM}}$ to align with the angular domain channel characteristics, the correlation between them can be maximized, leading to enhanced radiation efficiency.
-This formulation serves as the foundation for the following theoretical analysis and optimization.
+2, the angular space $\pi/2,\pi/2]$ is uniformly divided into $M$ sampling points, which can be expressed as $\vartheta_{m}=-\frac{\pi}{2}+\frac{\pi m}{M},~m=1,2,\ldots,M$.
+Specifically, each reconfigurable radiation pattern (illustrated by the red pattern in Fig.
+2) is represented across $M$ discrete angular directions by a vector
+$\mathbf{f}_{\mathrm{EM}}\in\mathbb{R}^{M},$ (1)
+with each element corresponding to radiation gain in a particular direction. Correspondingly, we introduce the virtual angular index channel for pattern-reconfigurable antennas, denoted as
+$\mathbf{h}_{\mathrm{EM}}\in\{0,1\}^{M},$ (2)
+which serves as an indexing vector indicating the spatial angular information of the user. Consequently, the effective EM domain gain can be mathematically expressed as
+$G_{\mathrm{EM}}=\mathbf{h}_{\mathrm{EM}}^{H}\mathbf{f}_{\mathrm{EM}}\in\mathbb{R},$ (3)
+which quantifies the radiation intensity toward the user’s actual spatial angle.
+By optimizing the radiation pattern $\mathbf{f}_{\mathrm{EM}}$ to closely align with the angular channel $\mathbf{h}_{\mathrm{EM}}$, we can maximize their correlation, thus significantly enhancing the overall radiation efficiency.
+This theoretical formulation establishes a foundation for the subsequent analysis and optimization.
 Figure 3: Tri-hybrid beamforming architecture.
-Inspired by the increased design DoF introduced by pattern-reconfigurable antennas, we integrate this advanced hardware component into the conventional hybrid beamforming architecture and propose a tri-hybrid beamforming framework that incorporates beamforming across three domains.
-Specifically, radiation beamforming in the EM domain is employed to dynamically shape the radiation pattern, analog beamforming in the RF domain is utilized to enhance array gain, and digital beamforming in the BB domain is implemented to mitigate multi-user interference.
+Inspired by the increased design DoF introduced by pattern-reconfigurable antennas, we propose integrating this advanced antenna technology into conventional hybrid beamforming architectures, and develop a novel tri-hybrid beamforming framework that simultaneously considers three distinct beamforming domains.
+Specifically, radiation beamforming within the EM domain dynamically shapes the antenna radiation patterns, analog beamforming in the RF domain enhances array gain, and digital beamforming in the BB domain mitigates multi-user interference.
+This comprehensive integration fully leverages the unique advantages of each domain, thereby achieving superior performance and efficiency in complex wireless communication scenarios.
 II-B Tri-hybrid Beamforming Architecture
@@
 3.
-Specifically, a base station (BS) equipped with a uniform linear array (ULA) serves multiple single-antenna user equipments (UEs).
-The BS employs reconfigurable antennas that dynamically adjust their radiation patterns, while each UE is equipped with a conventional omnidirectional antenna.
+Specifically, a base station (BS) equipped with a uniform linear array (ULA) serves $K$ single-antenna user equipments (UEs).
+The wireless channel between the BS and each UE comprises $L$ propagation path.
+The BS employs $N_{\mathrm{t}}$ reconfigurable antennas that dynamically adjust their radiation patterns, while each UE is equipped with a conventional omnidirectional antenna.
 The antenna spacing at the BS is set to $d={\lambda}/{2}$, where $\lambda$ is the center carrier wavelength.
-Let $\mathcal{N}_{\mathrm{t}}\triangleq\{1,2,\ldots,N_{\mathrm{t}}\}$ denote the set of BS antennas, $\mathcal{K}\triangleq\{1,2,\ldots,K\}$ denote the set of UEs, and $\mathcal{L}\triangleq\{1,2,\ldots,L\}$ denote the set of propagation paths between the BS and each UE.
 In the following, we mathematically formulate the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in the EM domain, the analog beamformer $\mathbf{F}_{\mathrm{RF}}$ in the RF domain and the digital beamformer $\mathbf{F}_{\mathrm{BB}}$ in the BB domain, establishing the foundation for the tri-hybrid beamforming architecture.
-EM domain: Building upon the radiation pattern model of a single pattern-reconfigurable antenna $\mathbf{f}_{\mathrm{EM}}$ presented in Section II-A, we extend this model to characterize the programmable radiation pattern of the entire antenna array at the BS, which consists of $N_{\mathrm{t}}$ reconfigurable antennas.
-The resulting radiation beamformer in the EM domain is expressed in the following matrix form
+EM domain: The radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ dynamically adjusts the amplitude responses of each reconfigurable antenna element, facilitating flexible control over energy distribution across diverse spatial directions.
+Extending the radiation pattern model $\mathbf{f}_{\mathrm{EM}}$ described in Section II-A for a single pattern-reconfigurable antenna, we characterize the collective programmable radiation pattern of the entire antenna array at the BS, comprising $N_{\mathrm{t}}$ reconfigurable antennas.
+Mathematically, the EM domain radiation beamformer is represented in the following matrix form
 $\mathbf{F}_{\mathrm{EM}}\triangleq\mathrm{blkdiag}\{\mathbf{f}_{\mathrm{EM},1},\mathbf{f}_{\mathrm{EM},2},\ldots,\mathbf{f}_{\mathrm{EM},N_{\mathrm{t}}}\}\in\mathbb{R}^{MN_{\mathrm{t}}\times N_{\mathrm{t}}},$ (4)
 where the $m$-th element of $\mathbf{f}_{\mathrm{EM},n}$ represents the radiation gain of the $n$-th antenna in the $m$-th angular direction.
-To ensure fairness in pattern design, the total energy of each antenna’s radiation pattern is restricted and normalized as $\|\mathbf{f}_{\mathrm{EM},n}\|^{2}=1,\forall n$.
-To further characterize the spatial mapping of the reconfigurable antennas, we introduce the virtual angular index channel matrix
-$\overline{\mathbf{H}}_{\mathrm{EM},k}\triangleq\mathbf{I}_{N_{\mathrm{t}}}\otimes\mathbf{H}_{\mathrm{EM},k}\in\{0,1\}^{MN_{\mathrm{t}}\times LN_{\mathrm{t}}},~{}\forall k,$ (2)
-where $\mathbf{H}_{\mathrm{EM},k}\triangleq[\mathbf{h}_{\mathrm{EM},k,1},\mathbf{h}_{\mathrm{EM},k,2},\ldots,\mathbf{h}_{\mathrm{EM},k,L}]\in\{0,1\}^{M\times L}$ that specifies the spatial angles associated with the $k$-th UE.
-Each column $\mathbf{h}_{\mathrm{EM},k,l}$ contains exactly one non-zero element, indicating the angular direction of the $l$-th propagation path for the $k$-th UE.
-Consequently, the radiation gain $\mathbf{G}_{\mathrm{EM},k}\in\mathbb{R}^{LN_{\mathrm{t}}\times N_{\mathrm{t}}}$ of all antennas towards the $k$-th UE can be formulated as
-$\mathbf{G}_{\mathrm{EM},k}$ $=\overline{\mathbf{H}}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}$ (3)
-$=\mathrm{blkdiag}\{\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{f}_{\mathrm{EM},1},\ldots,\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{f}_{\mathrm{EM},N_{\mathrm{t}}}\},$
-where the $l$-th entry of $\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{f}_{\mathrm{EM},n}$ represents the actual radiation gain of the $n$-th antenna towards the $l$-th direction for the $k$-th UE.
-In summary, the matrices $\mathbf{F}_{\mathrm{EM}}$ and $\overline{\mathbf{H}}_{\mathrm{EM},k}$ mathematically capture the additional degrees of freedom introduced by reconfigurable antennas, enabling a more flexible and adaptive radiation beamforming strategy in the EM domain.
-RF domain: We adopt a dynamic-subarray analog beamforming architecture, where each RF chain is adaptively connected to a disjoint subset of the total transmit antennas via a switch network.
-By fully exploiting the design flexibility of dynamic antenna selection, this architecture achieves an optimal balance between system efficiency and communication performance.
-Without loss of generality, we assume that the number of RF chains $N_{\mathrm{RF}}$ is not less than the number of UEs $K$, i.e., $N_{\mathrm{RF}}\geq K$, and the set of RF chains is denoted as $\mathcal{N}_{\mathrm{RF}}\triangleq\{1,2,\ldots,N_{\mathrm{RF}}\}$.
+To ensure fairness in pattern design, the total energy of each antenna’s radiation pattern is restricted and normalized as $\|\mathbf{f}_{\mathrm{EM},n}\|^{2}=1$.
+To further delineate the spatial characteristics of reconfigurable antennas, we introduce the virtual angular index channel matrix associated with the $l$-th propagation path for the $k$-th UE
+$\mathbf{H}_{\mathrm{EM},k,l}\triangleq\mathbf{I}_{N_{\mathrm{t}}}\otimes\mathbf{h}_{\mathrm{EM},k,l}\in\{0,1\}^{MN_{\mathrm{t}}\times N_{\mathrm{t}}},~\forall k,l,$ (5)
+where $\mathbf{h}_{\mathrm{EM},k,l}\in\{0,1\}^{M}$ explicitly specifies the spatial angles associated with the $k$-th UE.
+Each $\mathbf{h}_{\mathrm{EM},k,l}$ contains exactly one non-zero element, indicating the angular direction of the $l$-th propagation path for the $k$-th UE among $M$ angular sampling points.
+Consequently, the actual radiation gains of all antennas $\mathbf{G}_{\mathrm{EM},k,l}\in\mathbb{R}^{N_{\mathrm{t}}\times N_{\mathrm{t}}}$ can be formulated as
+$\mathbf{G}_{\mathrm{EM},k,l}$ $=\mathbf{H}_{\mathrm{EM},k,l}^{H}\mathbf{F}_{\mathrm{EM}}$ (6)
+$\triangleq\mathrm{diag}\{G_{\mathrm{EM},k,l,1},\ldots,G_{\mathrm{EM},k,l,N_{\mathrm{t}}}\},~\forall k,l,$
+which is a diagonal matrix with the $n$-th entry $G_{\mathrm{EM},k,l,n}\triangleq\mathbf{h}_{\mathrm{EM},k,l}^{H}\mathbf{f}_{\mathrm{EM},n}$ representing the actual radiation gain of the $n$-th antenna towards the $l$-th path of the $k$-th UE.
+In summary, the matrices $\mathbf{F}_{\mathrm{EM}}$ and $\mathbf{H}_{\mathrm{EM},k,l}$ mathematically capture the additional DoF afforded by reconfigurable antennas, enabling highly adaptive radiation beamforming in the EM domain.
+RF domain: The analog beamformer $\mathbf{F}_{\mathrm{RF}}$ controls the phases of the transmitted signals across antennas, ensuring coherent signal combination and directional beamforming toward desired UEs.
+We adopt a dynamic-subarray analog beamforming architecture, wherein each RF chain adaptively connects to distinct subsets of antennas through a switching network.
+This architecture optimally balances system efficiency and communication performance by exploiting dynamic antenna selection.
+Assuming the number of RF chains $N_{\mathrm{RF}}$ meets or exceeds the number of UEs $K$, i.e., $N_{\mathrm{RF}}\geq K$.
 The analog beamformer employs a set of low-resolution PSs, each following a constant modulus constraint with quantized phase control of $B$ bits.
@@
 Moreover, to ensure non-overlapping subarrays, each antenna is dynamically assigned to only one RF chain, meaning that the analog beamformer matrix is restricted to have only one non-zero element per row, i.e., $\|\mathbf{F}_{\mathrm{RF}}(n,:)\|_{0}=1$.
-BB domain: We exploit the digital beamformer $\mathbf{F}_{\mathrm{BB}}\triangleq[\mathbf{f}_{\mathrm{BB},1},\mathbf{f}_{\mathrm{BB},2},\ldots,\mathbf{f}_{\mathrm{BB},K}]\in\mathbb{C}^{N_{\mathrm{RF}}\times K}$ to pre-process transmitted signals in the BB domain, addressing the limitation of having fewer RF chains, which restricts direct control over each antenna’s signal.
-This procedure effectively mitigates multi-user interference and enhances overall system performance.
+BB domain: The digital beamformer $\mathbf{F}_{\mathrm{BB}}$ operates at the baseband level to effectively suppress multi-user interference and optimize overall system performance.
+Specifically, it is defined as $\mathbf{F}_{\mathrm{BB}}\triangleq[\mathbf{f}_{\mathrm{BB},1},\mathbf{f}_{\mathrm{BB},2},\ldots,\mathbf{f}_{\mathrm{BB},K}]\in\mathbb{C}^{N_{\mathrm{RF}}\times K}$.
+By performing precise pre-processing of transmitted signals, the digital beamformer compensates for the limited number of RF chains, ensuring refined control over each user’s signal in the spatial domain and significantly enhancing system throughput.
 II-C Channel Model
 After establishing the mathematical formulation of the tri-hybrid beamforming architecture, our focus now shifts to developing a comprehensive channel model that accommodates the expanded dimensionality introduced by the pattern-reconfigurable antenna array.
-Specifically, unlike conventional channel models that primarily account for path loss and spatial phase gradients, the proposed model incorporates the virtual angular index channel matrix $\overline{\mathbf{H}}_{\mathrm{EM},k},\forall k$ in (2).
-This integration enables an explicit representation of the angular characteristics of each propagation path.
-Furthermore, when combined with the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in (1), it enables the determination of the radiation intensity $\mathbf{G}_{\mathrm{EM},k},\forall k$ in (3), facilitating a more precise characterization of the system’s electromagnetic behavior.
-To establish a foundation for the innovative channel model of the pattern-reconfigurable antenna array, we first present the traditional mmWave channel model.
-Due to the sparse scattering characteristics of the mmWave band, the channel typically consists of a few dominant clusters, each containing multiple propagation paths [28].
-Let $\mathcal{C}\triangleq\{1,2,\ldots,C\}$ denote the set of scattering clusters, and $\mathcal{L}_{c}\triangleq\{1,2,\ldots,L_{c}\}$ denote the set of multi-path rays within the $c$-th scattering cluster.
+Specifically, unlike conventional channel models that primarily account for path loss and spatial phase gradients, our proposed model explicitly integrates the virtual angular index channel matrix $\mathbf{H}_{\mathrm{EM},k,l}$, defined in (5), to represent the angular characteristics of each propagation path.
+By incorporating this channel matrix along with the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ described in (4), our approach precisely determines the radiation gain $\mathbf{G}_{\mathrm{EM},k,l}$ in (6), enabling a more accurate and comprehensive characterization of the EM domain behavior.
+To establish a foundation for the innovative channel model of the pattern-reconfigurable antenna array, we first revisit the traditional mmWave channel model.
+Given the sparse scattering characteristics of the mmWave band, the channel typically consists of $C$ dominant scattering clusters, each containing $L_{c}$ multi-path rays [31], [32].
 The channel between the $n$-th antenna and the $k$-th UE can be expressed as
-$h_{k,n}=\sum_{c=1}^{C}\sum_{l_{c}=1}^{L_{c}}\alpha_{k,c,l_{c}}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,c,l_{c}}},~{}\forall k,n,$ (4)
-where $\theta_{k,c,l_{c}}$ and $\alpha_{k,c,l_{c}}$ denote the angle of departure (AoD) and channel gain corresponding to the $l_{c}$-th ray in the $c$-th scattering cluster, respectively.
+$h_{k,n}=\sum_{c=1}^{C}\sum_{l_{c}=1}^{L_{c}}\alpha_{k,c}e^{j\phi_{k,l_{c}}}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,c,l_{c}}},~\forall k,n,$ (7)
+where $\theta_{k,c,l_{c}}$ denotes the angle of departure (AoD) corresponding to the $l_{c}$-th ray in the $c$-th scattering cluster, respectively, $\alpha_{k,c}$ denotes the mean large-scale channel gain of the $c$-th scattering cluster, and $\phi_{k,l_{c}}$ denotes the small-scale phase term of the $l_{c}$-th ray.
 The AoD $\theta_{k,c,l_{c}}$ can be described as
@@
 where $\overline{\theta}_{k,c}$ denotes the nominal AoD of the $c$-th scattering cluster for the $k$-th UE, and $\Delta\theta_{k,l_{c}}$ denotes the angular deviation of the $l_{c}$-th ray from the nominal direction, following Gaussian distribution $\Delta\theta_{k,l_{c}}\sim\mathcal{N}(0,\varsigma^{2}_{k,c})$, with $\varsigma_{k,c}$ indicating the angle spread.
-The channel gain involves the typically distance-dependent path loss with $r_{k,c,l_{c}}$ representing the propagation distance.
-It follows the same modeling as in (5), i.e., $r_{k,c,l_{c}}=\overline{r}_{k,c}+\Delta r_{k,l_{c}}$, where $\Delta r_{k,l_{c}}\sim\mathcal{N}(0,\iota^{2}_{k,c})$.
+The mean large-scale channel gain $\alpha_{k,c}$ mainly involves the typically distance-dependent path loss with $r_{k,c}$ representing the propagation distance.
+The small-scale parameter $\phi_{k,l_{c}}$ follows a uniform distribution in $0,2\pi]$ and is independent from ray to ray.
 For better clarity in the subsequent channel model formulation, we define $L\triangleq\sum_{c=1}^{C}L_{c}$ as the total number of paths.
-With this notation, the channel model in (4) can be reformulated in a more compact form as follows
-$h_{k,n}=\sum_{l=1}^{L}\alpha_{k,l}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,l}},~{}\forall k,n.$ (6)
-Furthermore, the integration of pattern-reconfigurable antennas extends the channel representation from a one-dimensional to an $M$-dimensional formulation by incorporating additional EM domain channel information, i.e., $\overline{\mathbf{H}}_{\mathrm{EM},k},\forall k$ in (2).
-Specifically, the dimension-extended channel for the $n$-th pattern-reconfigurable antenna, denoted as $\mathbf{h}_{k,n}\in\mathbb{C}^{M}$, is expressed as follows
-$\mathbf{h}_{k,n}=\sum_{l=1}^{L}(\alpha_{k,l}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,l}})\otimes\mathbf{h}_{\mathrm{EM},k,l},~{}\forall k,n.$ (7)
-Following this formulation, by stacking the channel vectors $\mathbf{h}_{k,n},\forall n$ of all $N_{\mathrm{t}}$ antennas, we obtain the complete channel vector for the $k$-th UE, given by $\mathbf{h}_{k}\triangleq[\mathbf{h}_{k,1}^{T},\mathbf{h}_{k,2}^{T},\ldots,\mathbf{h}_{k,N_{\mathrm{t}}}^{T}]^{T}\in\mathbb{C}^{MN_{\mathrm{t}}}$.
+With this notation, the channel model in (7) can be reformulated in a more compact form as
+$h_{k,n}=\sum_{l=1}^{L}\alpha_{k,l}e^{\jmath\phi_{k,l}}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,l}},~\forall k,n.$ (9)
+Furthermore, the integration of pattern-reconfigurable antennas extends the channel representation from a one-dimensional to an $M$-dimensional formulation by incorporating additional EM domain channel information, i.e., $\mathbf{H}_{\mathrm{EM},k,l}$ in (5).
+Specifically, the dimension-extended channel for the $n$-th pattern-reconfigurable antenna $\mathbf{h}_{k,n}\in\mathbb{C}^{M}$ is expressed as
+$\mathbf{h}_{k,n}\!\!=\!\!\sum_{l=1}^{L}(\alpha_{k,l}e^{\jmath\phi_{k,l}}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,l}}\!)\!\otimes\!\mathbf{h}_{\mathrm{EM},k,l},~\forall k,n.\!\!\!\!$ (10)
+Following this formulation, by stacking the channel vectors $\mathbf{h}_{k,n}$ of all antennas, we obtain the comprehensive channel vector for the $k$-th UE, given by $\mathbf{h}_{k}\triangleq[\mathbf{h}_{k,1}^{T},\mathbf{h}_{k,2}^{T},\ldots,\mathbf{h}_{k,N_{\mathrm{t}}}^{T}]^{T}\in\mathbb{C}^{MN_{\mathrm{t}}}$.
 The corresponding expression can be explicitly formulated as
-$\mathbf{h}_{k}=\sum_{l=1}^{L}(\alpha_{k,l}\mathbf{a}_{k,l})\otimes\mathbf{h}_{\mathrm{EM},k,l},~{}\forall k,$ (8)
-where
-$\mathbf{a}_{k,l}$ $\triangleq[1,e^{-\jmath\frac{2\pi}{\lambda}d\sin\theta_{k,l}},\ldots,e^{-\jmath\frac{2\pi}{\lambda}(N_{\mathrm{t}}-1)d\sin\theta_{k,l}}]^{T},~{}\forall k,l,$ (9)
-denotes the steering vector. By applying the Kronecker product formula and rearranging the order of summation, (8) can be equivalently expressed as
-$\mathbf{h}_{k}=\mathrm{vec}(\mathbf{H}_{\mathrm{EM},k}\mathbf{H}_{\mathrm{S},k}^{T}),~{}\forall k,$ (10)
-with
-$\mathbf{H}_{\mathrm{S},k}$ $=[\alpha_{k,1}\mathbf{a}_{k,1},\alpha_{k,2}\mathbf{a}_{k,2},\ldots,\alpha_{k,L}\mathbf{a}_{k,L}]\in\mathbb{C}^{N_{\mathrm{t}}\times L},~{}\forall k,$ (11)
-representing the spatial domain channel.
-Additionally, to enhance clarity and brevity, the EM domain channel and spatial domain channel can be separated into two independent multiplicative expressions, which yields
-$\mathbf{h}_{k}=\overline{\mathbf{H}}_{\mathrm{EM},k}\overline{\mathbf{h}}_{\mathrm{S},k},~{}\forall k,$ (12)
-where $\overline{\mathbf{H}}_{\mathrm{EM},k}\in\{0,1\}^{MN_{\mathrm{t}}\times LN_{\mathrm{t}}}$ is defined in (2) containing the EM domain channel information, and $\overline{\mathbf{h}}_{\mathrm{S},k}\triangleq\mathrm{vec}(\mathbf{H}_{\mathrm{S},k}^{T})\in\mathbb{C}^{LN_{\mathrm{t}}}$ includes the spatial domain channel information.
+$\mathbf{h}_{k}$ $=\sum_{l=1}^{L}(\alpha_{k,l}e^{\jmath\phi_{k,l}}\mathbf{a}_{k,l})\otimes\mathbf{h}_{\mathrm{EM},k,l},$ (11)
+$\triangleq\sum_{l=1}^{L}\underbrace{(\mathbf{I}_{N_{\mathrm{t}}}\otimes\mathbf{h}_{\mathrm{EM},k,l})}_{\mathbf{H}_{\mathrm{EM},k,l}}\underbrace{(\alpha_{k,l}e^{\jmath\phi_{k,l}}\mathbf{a}_{k,l})}_{\mathbf{h}_{\mathrm{S},k,l}},~\forall k,$
+where the steering vector is defined as
+$\!\!\!\mathbf{a}_{k,l}$ $\!\triangleq\![1,e^{-\jmath\frac{2\pi}{\lambda}d\sin\theta_{k,l}},\ldots,e^{-\jmath\frac{2\pi}{\lambda}(N_{\mathrm{t}}-1)d\sin\theta_{k,l}}]^{T},~\forall k,l.\!\!\!\!\!\!\!\!$ (12)
+The comprehensive channel $\mathbf{h}_{k}$ in (11) is composed of the virtual angular index channel matrix $\mathbf{H}_{\mathrm{EM},k,l}$ defined in (5) and the spatial domain channel $\mathbf{h}_{\mathrm{S},k,l}$ given by
+$\mathbf{h}_{\mathrm{S},k,l}\triangleq\alpha_{k,l}e^{\jmath\phi_{k,l}}\mathbf{a}_{k,l},~\forall k,l.$ (13)
+When the virtual angular index channel matrix $\mathbf{H}_{\mathrm{EM},k,l}$ interacts with the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in (4), the actual radiation gains of the antennas $\mathbf{G}_{\mathrm{EM},k,l}$ mentioned in (6) are obtained.
+These gains subsequently modify the conventional spatial domain channel $\mathbf{h}_{\mathrm{S},k,l}$, thereby shaping the effective channel seen by the analog and digital beamformers.
+Additionally, to enhance clarity and brevity, we replace the path-wise summation in (11) with an equivalent block-stacked multiplicative expressions, which yields
+$\mathbf{h}_{k}\triangleq\mathbf{H}_{\mathrm{EM},k}\mathbf{h}_{\mathrm{S},k},~\forall k,$ (14)
+where $\mathbf{H}_{\mathrm{EM},k}\triangleq[\mathbf{H}_{\mathrm{EM},k,1},\mathbf{H}_{\mathrm{EM},k,2},\ldots,\mathbf{H}_{\mathrm{EM},k,L}]\in\{0,1\}^{MN_{\mathrm{t}}\times LN_{\mathrm{t}}}$ contains the virtual angular index channel matrices, and $\mathbf{h}_{\mathrm{S},k}\triangleq[\mathbf{h}_{\mathrm{S},k,1}^{T},\mathbf{h}_{\mathrm{S},k,2}^{T},\ldots,\mathbf{h}_{\mathrm{S},k,L}^{T}]^{T}\in\mathbb{C}^{LN_{\mathrm{t}}}$ includes the spatial domain channel information.
 II-D System Model
@@
 Based on the above mathematical definition, the signal received by the $k$-th UE is given by
-$y_{k}$ $=\overline{\mathbf{h}}_{\mathrm{S},k}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}\mathbf{x}+n_{k},~{}\forall k,$ (13)
+$y_{k}$ $=\mathbf{h}_{k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}\mathbf{x}+n_{k}$ (15)
+$=\mathbf{h}_{\mathrm{S},k}^{H}\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}\mathbf{x}+n_{k},~\forall k,$
 where $\mathbf{x}\triangleq[x_{1},x_{2},\ldots,x_{K}]^{T}\in\mathbb{C}^{K}$ denotes the transmitted symbols of $K$ UEs that satisfies $\mathbb{E}\{\mathbf{x}\mathbf{x}^{H}\}=\mathbf{I}_{K}$, and $n_{k}\!\sim\!\mathcal{CN}(0,\sigma^{2}_{k})$ denotes the additive white Gaussian noise (AWGN) at the $k$-th UE.
 The signal-to-interference-plus-noise ratio (SINR) of the $k$-th UE can be calculated as
-$\mathrm{SINR}_{k}=\frac{|\mathbf{\overline{h}}_{\mathrm{S},k}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},k}|^{2}}{\sum_{j=1,j\neq k}^{K}|\mathbf{\overline{h}}_{\mathrm{S},k}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},j}|^{2}+\sigma_{k}^{2}},~{}\forall k,\$ (14)
+$\mathrm{SINR}_{k}=\frac{|\mathbf{h}_{\mathrm{S},k}^{H}\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},k}|^{2}}{\sum_{j=1,j\neq k}^{K}|\mathbf{h}_{\mathrm{S},k}^{H}\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},j}|^{2}+\sigma_{k}^{2}},~\forall k,\\$ (16)
 and the corresponding achievable rate can be calculated as
@@
 II-E Problem Formulation
-In this paper, based on the tri-hybrid beamforming architecture, we propose to jointly design the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in the EM domain, the analog beamformer $\mathbf{F}_{\mathrm{RF}}$ in the RF domain, and the digital beamformer $\mathbf{F}_{\mathrm{BB}}$ in the BB domain to maximize the sum-rate, while satisfying the constraints of radiation pattern energy, constant modulus and discrete phases of PSs, non-overlapping dynamic connection network, and total transmit power budget.
-Thus, the joint three domain beamformer optimization problem with full instantaneous CSI is formulated as
+In this paper, we propose a comprehensive tri-hybrid beamforming strategy, jointly optimizing the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in the EM domain, the analog beamformer $\mathbf{F}_{\mathrm{RF}}$ in the RF domain, and the digital beamformer $\mathbf{F}_{\mathrm{BB}}$ in the BB domain.
+The primary objective is to maximize the sum-rate performance of the system while satisfying several critical constraints, including radiation pattern energy normalization, constant modulus and discrete-phase constraints of the PSs, non-overlapping dynamic subarray assignments, and a total transmit power budget.
+Therefore, the joint optimization problem encompassing the three beamforming domains under the assumption of perfect instantaneous CSI is formulated as
 $\max_{\mathbf{F}_{\mathrm{EM}},\mathbf{F}_{\mathrm{RF}},\mathbf{F}_{\mathrm{BB}}}$ $\ \ \sum^{K}_{k=1}R_{k}$ (18a)
-$\mathrm{s.t.}$ $\ \ \|\mathbf{f}_{\mathrm{EM},n}\|^{2}=1,~{}\forall n,$ (16b)
+$\mathrm{s.t.}\ \ \ \ \$$\ \ \|\mathbf{f}_{\mathrm{EM},n}\|^{2}=1,~\forall n,$ (18b)
 $\ \ \mathbf{F}_{\mathrm{RF}}(n,l_{\mathrm{RF}})\in\{\mathcal{F},0\},~\forall n,l_{\mathrm{RF}},$ (18c)
@@
 where $P_{\mathrm{t}}$ is the transmit power budget.
-Within this tri-hybrid beamforming architecture, accurately acquiring instantaneous CSI is a significant challenge.
-The analog beamformer in the RF domain suffers from inherent undersampling, while the need to estimate instantaneous CSI for all possible radiation patterns in the EM domain leads to excessive pilot overhead.
-Moreover, the real-time computation of beamforming matrices across all three domains imposes high computational complexity, severely impacting system processing efficiency.
-To address these issues, we propose an innovative tri-timescale framework tailored for tri-hybrid beamforming design.
-This framework decomposes the original non-convex optimization problem in (16) into three independent subproblems, each solvable at a different timescale.
-This decomposition significantly reduces pilot overhead, enhances computational tractability, and improves overall system efficiency.
+Accurately obtaining instantaneous CSI poses significant practical difficulties within this tri-hybrid framework.
+In the RF domain, the analog beamformer inherently suffers from undersampling due to the limited number of RF chains, while acquiring instantaneous CSI for all possible radiation patterns in the EM domain leads to excessive pilot overhead.
+In addition, real-time joint optimization of the beamforming matrices across all three domains results in high computational complexity and low processing efficiency.
+These challenges motivate the development of a more structured beamforming strategy, which will be addressed in the next section.
+Figure 4: The frame structure of tri-timescale tri-hybrid beamforming.
 III Tri-timescale Beamforming Designs
 III-A Tri-timescale Framework
-The tri-hybrid beamforming architecture presents significant practical challenges, including high channel estimation overhead and computational complexity.
-Estimating the full CSI and jointly optimizing beamformers across all three domains in every time-slot is computationally infeasible.
-To address this, we propose a tri-timescale beamforming framework that effectively balances practical feasibility and system performance, as illustrated in Fig.
+In this section, we present a tri-timescale beamforming framework for the proposed tri-hybrid architecture.
+Instead of estimating full instantaneous CSI and jointly optimizing the radiation, analog, and digital beamformers in every time-slot, the design separates their updates across three different timescales so as to balance performance and implementation complexity, as illustrated in Fig.
 4.
-The overall time domain corresponds to the channel statistics coherence time, during which the statistical distribution of the channel remains stable.
-This period is structured hierarchically, with beamforming operations executed across three distinct timescales, as detailed below:
-  • Long-timescale: The total duration is divided into $T_{\mathrm{L}}$ super-frames, denoted as $\mathcal{T}_{\mathrm{L}}\triangleq\{1,2,\ldots,T_{\mathrm{L}}\}$.
+The overall time horizon corresponds to the channel statistics coherence time, during which the statistical distribution of the channel remains approximately constant [^1] Our channel modeling follows the quasi-static clustered model described in 3GPP TR 38.901 [31], where cluster-level parameters (e.g., nominal angles and angle spreads) remain approximately constant within a channel-statistics coherence interval.
+This assumption is well suited for low-to-moderate mobility scenarios such as indoor hotspots and small-cell mmWave deployments.
+Extending the proposed tri-timescale framework to fully dynamic, high-mobility scenarios is an important topic for future research..
+This period is organized hierarchically, with beamforming operations carried out over three distinct timescales, as detailed below:
+  • Long-timescale: The total duration, corresponding to the channel-statistics coherence time, is divided into $T_{\mathrm{L}}$ super-frames, denoted as $\mathcal{T}_{\mathrm{L}}\triangleq\{1,2,\ldots,T_{\mathrm{L}}\}$.
+This coherence interval typically spans from a few seconds to several tens of seconds.
 In each super-frame, the BS extracts high-dimensional channel information of the scattering cluster core to design the radiation beamformer, ensuring radiation pattern alignment in the EM domain.
-  • Medium-timescale: Each super-frame consists of $T_{\mathrm{M}}$ frames, defined as $\mathcal{T}_{\mathrm{M}}\triangleq\{1,2,\ldots,T_{\mathrm{M}}\}$.
+  • Medium-timescale: Each super-frame consists of $T_{\mathrm{M}}$ frames, denoted as $\mathcal{T}_{\mathrm{M}}\triangleq\{1,2,\ldots,T_{\mathrm{M}}\}$.
+The duration of each frame is determined by the sampling frequency of the statistical channel information, generally ranging from several hundred milliseconds to a few seconds.
 In each frame, the BS acquires a statistical channel sample of multi-path rays within the scattering clusters to update the analog beamformer, enhancing array gain in the RF domain.
-  • Short-timescale: Each frame is further divided into $T_{\mathrm{S}}$ time-slots, referred to as the channel coherence time, represented as $\mathcal{T}_{\mathrm{S}}\triangleq\{1,2,\ldots,T_{\mathrm{S}}\}$.
+  • Short-timescale: Each frame is further divided into $T_{\mathrm{S}}$ time-slots, denoted as $\mathcal{T}_{\mathrm{S}}\triangleq\{1,2,\ldots,T_{\mathrm{S}}\}$.
+Each time-slot is referred to as the channel coherence time, typically on the order of tens of milliseconds [^2] The durations of the three timescales depend on the carrier frequency, user mobility, and propagation environment.
+The orders of magnitude reported here follow typical mmWave deployments under the 3GPP TR 38.901 clustered channel model and serve only as representative examples [29].
+In practice, the time-scale durations can be flexibly adjusted to match specific deployment scenarios without altering the proposed tri-timescale optimization framework..
 In each time-slot, the BS obtains an instantaneous low-dimensional effective CSI to refine the digital beamformer, mitigating multi-user interference and accounting for fast channel fluctuations in the BB domain.
@@
 In the following subsections, we provide a detailed description of the channel estimation processes and beamforming design procedures based on this tri-timescale framework.
-Figure 4: The frame structure of tri-timescale tri-hybrid beamforming.
 III-B Long-timescale EM Domain Design
@@
 Instead, aligning radiation patterns with the nominal angles of scattering clusters is sufficient to maintain robust communication performance, while significantly reducing channel estimation overhead and beamforming design complexity.
-Therefore, over the long-timescale, our objective is to estimate the nominal angles and average channel gains of the scattering clusters, which serve as the foundation for radiation beamforming optimization.
-Specifically, we employ the Capon algorithm to construct the angular power spectrum (APS) based on the channel covariance matrix, enabling precise spatial domain analysis [29].
+Therefore, over the long-timescale, our objective is to estimate the nominal angles $\overline{\theta}_{k,c}$ and the mean large-scale channel gains $\alpha_{k,c}$ of the scattering clusters, which serve as the foundation for radiation beamforming optimization.
+Specifically, we employ the Capon algorithm to construct the angular power spectrum (APS) based on the channel covariance matrix, enabling the extraction of scattering cluster channel parameters [32], [33].
 To illustrate the channel estimation process, we consider the $k$-th UE as an example.
-By accumulating spatial-channel samples over the medium-timescale, the average channel covariance matrix at the $t_{\mathrm{L}}$-th super-frame is given by
-$\mathbf{R}_{k}^{(t_{\mathrm{L}})}=\frac{1}{t_{\mathrm{L}}T_{\mathrm{M}}}\sum_{t_{\mathrm{l}}=1}^{t_{\mathrm{L}}}\sum_{t_{\mathrm{M}}=1}^{T_{\mathrm{M}}}\widetilde{\mathbf{h}}^{(t)}_{\mathrm{S},k}(\widetilde{\mathbf{h}}_{\mathrm{S},k}^{(t)})^{H},~{}\forall k,t_{\mathrm{L}},$ (17)
-where $t=(t_{\mathrm{l}}-1)T_{\mathrm{M}}+t_{\mathrm{M}}$ is defined for brevity and $\widetilde{\mathbf{h}}_{\mathrm{S},k}$ represents the estimated channel excluding the influence of the radiation patterns, given by
-$\widetilde{\mathbf{h}}_{\mathrm{S},k}=\sum_{l=1}^{L}\alpha_{k,l}\mathbf{a}_{k,l},~{}\forall k.$ (18)
+In the first super-frame, all the pattern-reconfigurable antennas are typically configured to radiate toward the boresight direction.
+Starting from the second super-frame, the scattering-cluster channel information is estimated by accumulating spatial channel samples collected over the medium-timescale of the preceding super-frames.
+Accordingly, the channel covariance matrix at the $t_{\mathrm{L}}$-th super-frame is given by
+$\mathbf{R}_{k}^{(t_{\mathrm{L}})}=\frac{1}{t_{\mathrm{L}}T_{\mathrm{M}}}\sum_{t_{\mathrm{l}}=1}^{t_{\mathrm{L}}-1}\sum_{t_{\mathrm{M}}=1}^{T_{\mathrm{M}}}\widetilde{\mathbf{h}}^{(t)}_{\mathrm{S},k}(\widetilde{\mathbf{h}}_{\mathrm{S},k}^{(t)})^{H},~\forall k,t_{\mathrm{L}},$ (19)
+where $t=(t_{\mathrm{l}}-1)T_{\mathrm{M}}+t_{\mathrm{M}}$ is defined for brevity and $\widetilde{\mathbf{h}}_{\mathrm{S},k}\triangleq\sum_{l}^{L}\alpha_{k,l}e^{j\phi_{k,l}}\mathbf{a}_{k,l}$ represents the estimated channel excluding the influence of the radiation patterns.
 Then, the APS can be derived as
-$\rho^{(t_{\mathrm{L}})}_{k}(\theta)=\frac{1}{\mathbf{e}^{H}(\theta)(\mathbf{R}_{k}^{(t_{\mathrm{L}})})^{-1}\mathbf{e}(\theta)},~{}\forall k,t_{\mathrm{L}},$ (19)
-with
-$\mathbf{e}(\theta)\triangleq[1,e^{-\jmath\frac{2\pi}{\lambda}d\sin\theta},\ldots,e^{-\jmath\frac{2\pi}{\lambda}(N_{\mathrm{t}}-1)d\sin\theta}]^{T}\in\mathbb{C}^{N_{\mathrm{t}}}.$ (20)
-Since the APS reflects the power distribution across the angular domain, each peak in the APS physically represents the presence of a distinct scattering cluster in the propagation environment.
-Thus, the number of the peaks in $\rho^{(t_{\mathrm{L}})}_{k}(\theta)$ that exceed a specified threshold directly provides the estimated number of scattering clusters $C$.
-For each of these peaks, the corresponding angle value and power value indicate the estimated nominal angle $\overline{\theta}_{k,c}$ and the average channel gain $\alpha_{k,c}$ of the corresponding scattering cluster, respectively.
-In conclusion, referring to (12), we can construct the EM domain channel, which serves as the foundation for the long-timescale EM domain radiation beamformer design, which yields
-$\mathbf{h}_{\mathrm{C},k}=\overline{\mathbf{H}}_{\mathrm{EM},k}\overline{\mathbf{h}}_{\mathrm{S},k}\in\mathbb{C}^{MN_{\mathrm{t}}},$ (21)
-where the term $\overline{\mathbf{h}}_{\mathrm{S},k}\in\mathbb{C}^{CN_{\mathrm{t}}}$ encapsulates the spatial domain channel characteristics of $C$ scattering cluster cores and $\mathbf{\overline{H}}_{\mathrm{EM},k}\in\{0,1\}^{MN_{\mathrm{t}}\times CN_{\mathrm{t}}}$ indicates the angle information of the scattering clusters.
+$\rho^{(t_{\mathrm{L}})}_{k}(\vartheta_{m})=\frac{1}{\mathbf{e}^{H}(\vartheta_{m})(\mathbf{R}_{k}^{(t_{\mathrm{L}})})^{-1}\mathbf{e}(\vartheta_{m})},~\forall k,t_{\mathrm{L}},$ (20)
+with $\mathbf{e}(\vartheta_{m})\!\triangleq\![1,e^{-\jmath\frac{2\pi}{\lambda}d\sin\vartheta_{m}},\ldots,e^{-\jmath\frac{2\pi}{\lambda}(N_{\mathrm{t}}-1)d\sin\vartheta_{m}}]^{T}\in\mathbb{C}^{N_{\mathrm{t}}}.$ Since the APS reflects the power distribution across the angular domain, each peak in the APS physically represents the presence of a distinct scattering cluster in the propagation environment.
+Thus, the number of the peaks in $\rho^{(t_{\mathrm{L}})}_{k}(\vartheta_{m})$ that exceed a specified threshold directly provides the estimated number of scattering clusters $C$.
+For each of these peaks, the corresponding angle value and power value indicate the estimated nominal angle $\overline{\theta}_{k,c}$ and the mean large-scale channel gain $\alpha_{k,c}$ of the corresponding scattering cluster, respectively.
+In conclusion, referring to (14), we can construct the channels of scattering clusters as
+$\mathbf{h}_{\mathrm{C},k}=\mathbf{H}_{\mathrm{EM},\mathrm{C},k}\mathbf{h}_{\mathrm{S},\mathrm{C},k}\in\mathbb{C}^{MN_{\mathrm{t}}},~\forall k,$ (21)
+where $\mathbf{H}_{\mathrm{EM},\mathrm{C},k}\triangleq[\mathbf{H}_{\mathrm{EM},\mathrm{C},k,1},\mathbf{H}_{\mathrm{EM},\mathrm{C},k,2},\ldots,\mathbf{H}_{\mathrm{EM},\mathrm{C},k,C}]\in\{0,1\}^{MN_{\mathrm{t}}\times CN_{\mathrm{t}}}$ indicates the EM domain angle information and $\mathbf{h}_{\mathrm{S},\mathrm{C},k}\triangleq[\mathbf{h}_{\mathrm{S},\mathrm{C},k,1}^{T},\mathbf{h}_{\mathrm{S},\mathrm{C},k,2}^{T},\ldots,\mathbf{h}_{\mathrm{S},\mathrm{C},k,C}^{T}]^{T}\in\mathbb{C}^{CN_{\mathrm{t}}}$ encapsulates the spatial domain channel characteristics of $C$ scattering cluster cores.
+It serve as the foundation for the long-timescale EM domain radiation beamformer design.
 III-B.2 Long-timescale Radiation Beamformer Design
@@
 This optimization requires only the channel information of the scattering cluster core, i.e., the estimated nominal angles and average channel gains of the scattering clusters.
-Consequently, we substitute the instantaneous multi-path channel $\mathbf{h}_{k}$ in the SINR expression (14) and the sum-rate formulation (15) with the scattering cluster core channel $\mathbf{h}_{\mathrm{C},k}$, defining the long-timescale sum-rate $\overline{R}_{k}$ as our optimization objective.
+Consequently, we substitute the instantaneous multi-path channel $\mathbf{h}_{k}$ in the SINR expression (16) and the sum-rate formulation (17) with the scattering cluster core channel $\mathbf{h}_{\mathrm{C},k}$, and define the long-timescale sum-rate $R_{\mathrm{C},k}$ as our optimization objective.
 Given a fixed analog beamformer $\mathbf{F}_{\mathrm{RF}}$ and digital beamformer $\mathbf{F}_{\mathrm{BB}}$, the long-timescale EM domain radiation beamforming design problem is then formulated as
-$\max_{\mathbf{F}_{\mathrm{EM}}}$ $\ \ \sum^{K}_{k=1}\overline{R}_{k}$ (22a)
+$\max_{\mathbf{F}_{\mathrm{EM}}}$ $\ \ \sum^{K}_{k=1}R_{\mathrm{C},k}$ (22a)
 $\mathrm{s.t.}$ $\ \ \|\mathbf{f}_{\mathrm{EM},n}\|^{2}=1,~\forall n.$ (22b)
@@
 where $\overline{\mathbf{F}}_{\mathrm{pat}}\triangleq\mathbf{I}_{N_{\mathrm{t}}}\otimes\mathbf{F}_{\mathrm{pat}}\in\mathbb{R}^{MN_{\mathrm{t}}\times PN_{\mathrm{t}}}$ with $\mathbf{F}_{\mathrm{pat}}\in\mathbb{R}^{M\times P}$ representing a predefined radiation pattern dictionary that contains the radiation gains of $P$ candidate patterns over $M$ spatial sampling angles.
-The matrix $\mathbf{S}_{\mathrm{EM}}\triangleq\mathrm{blkdiag}\{\mathbf{s}_{\mathrm{EM},1},\mathbf{s}_{\mathrm{EM},2},\ldots,\mathbf{s}_{\mathrm{EM},N_{\mathrm{t}}}\}\in\{0,1\}^{PN_{\mathrm{t}}\times N_{\mathrm{t}}}$ denotes the EM domain pattern selection matrix, where each $\mathbf{s}_{\mathrm{EM},n}\in\{0,1\}^{P}$ selects a single optimal pattern from the dictionary for the $n$-th antenna, i.e., $\|\mathbf{s}_{\mathrm{EM},n}\|_{1}=1$.
-Based on this formulation, the original problem in (22) is equivalently transformed into a more practical optimization problem focused on optimizing the EM-domain pattern selection matrix $\mathbf{S}_{\mathrm{EM}}$, given by
-$\max_{\mathbf{S}_{\mathrm{EM}}}$ $\ \ \sum^{K}_{k=1}\overline{R}_{k}$ (24a)
+The matrix $\mathbf{S}_{\mathrm{EM}}\triangleq\mathrm{blkdiag}\{\mathbf{s}_{\mathrm{EM},1},\mathbf{s}_{\mathrm{EM},2},\ldots,\mathbf{s}_{\mathrm{EM},N_{\mathrm{t}}}\}\in\{0,1\}^{PN_{\mathrm{t}}\times N_{\mathrm{t}}}$ denotes the EM domain pattern selection matrix, where each sub-vector $\mathbf{s}_{\mathrm{EM},n}\in\{0,1\}^{P}$ selects a single optimal pattern from the dictionary for the $n$-th antenna, i.e., $\|\mathbf{s}_{\mathrm{EM},n}\|_{1}=1$.
+Based on this formulation, the original problem in (22) is equivalently transformed into a more practical optimization problem focused on optimizing the EM domain pattern selection matrix $\mathbf{S}_{\mathrm{EM}}$, given by
+$\max_{\mathbf{S}_{\mathrm{EM}}}$ $\ \ \sum^{K}_{k=1}R_{\mathrm{C},k}$ (24a)
 $\mathrm{s.t.}$ $\ \ \mathbf{s}_{\mathrm{EM},n}(p)\in\{0,1\},~\forall n,p,$ (24b)
@@
 This enhancement not only improves channel estimation accuracy but also facilitates a more robust and adaptive beamforming strategy, ultimately enhancing overall system performance.
-Initially, the RF domain channel to be estimated is expressed as
-$\mathbf{h}_{\mathrm{RF},k}\triangleq\mathbf{F}_{\mathrm{EM}}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}\overline{\mathbf{h}}_{\mathrm{S},k},~{}\forall k,$ (37)
-where $\mathbf{F}_{\mathrm{EM}}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}$ reflects the radiation gains of the reconfigurable antennas based on the fixed radiation beamformer determined in the long-timescale, and it is apparent that this term is coupled with the spatial channel $\overline{\mathbf{h}}_{\mathrm{S},k}$.
-During the training phase, all UEs simultaneously transmit pilot signals on distinct frequency tones to the BS in the uplink.
-Taking the $k$-th UE as an example, the received signal at the BS is given by
-$\mathbf{y}_{k}$ $=\sqrt{P_{\mathrm{t}}}\mathbf{W}_{k}^{H}\mathbf{h}_{\mathrm{RF},k}+\mathbf{W}_{k}^{H}\mathbf{n}_{k}$ (38)
-$=\sqrt{P_{\mathrm{t}}}\mathbf{W}_{k}^{H}\mathbf{A}_{\mathrm{D}}\mathbf{z}_{k}+\mathbf{W}_{k}^{H}\mathbf{n}_{k}$
-$=\mathbf{Q}_{k}\mathbf{z}_{k}+\mathbf{W}_{k}^{H}\mathbf{n}_{k},~{}\forall k,$
-where $\mathbf{W}_{k}\in\mathbb{C}^{N_{\mathrm{t}}\times N_{\mathrm{Beam}}}$ is the training beamformer with $N_{\mathrm{Beam}}$ denoting the number of training beams, $\mathbf{A}_{\mathrm{D}}\in\mathbb{C}^{N_{\mathrm{t}}\times M}$ is the refined dictionary matrix incorporating the effects of pattern-reconfigurable antennas, $\mathbf{z}_{k}$ is a sparse vector with non-zero elements corresponding to the spatial angles, and $\mathbf{Q}_{k}\triangleq\sqrt{P_{\mathrm{t}}}\mathbf{W}_{k}^{H}\mathbf{A}_{\mathrm{D}}$ is the sensing matrix.
+Initially, the RF domain channel at each frame to be estimated is expressed as
+$\mathbf{h}_{\mathrm{RF},k}\triangleq\mathbf{F}_{\mathrm{EM}}^{H}\mathbf{H}_{\mathrm{EM},k}\mathbf{h}_{\mathrm{S},k},~\forall k,$ (37)
+where $\mathbf{F}_{\mathrm{EM}}^{H}\mathbf{H}_{\mathrm{EM},k}$ reflects the radiation gains of the reconfigurable antennas based on the fixed radiation beamformer determined in the long-timescale, and it is apparent that this term is coupled with the spatial channel $\mathbf{h}_{\mathrm{S},k}$.
+During the uplink training phase, we assume the $K$ UEs transmit pilot sequences over orthogonal time resources.
+Therefore, the BS can isolate each UE’s signal, making channel estimation for different UEs completely independent.
+Suppose that $I$ pilot symbols are allocated to each UE.
+Taking the $k$-th UE as an example, the corresponding received signal during the $i$-th pilot at the BS $\mathbf{y}_{k}[i]\in\mathbb{C}^{N_{\mathrm{RF}}}$ can be expressed as
+$\mathbf{y}_{k}[i]$ $=\mathbf{W}[i]^{H}\mathbf{h}_{\mathrm{RF},k}x[i]+\mathbf{W}[i]^{H}\mathbf{n}[i],~\forall k,i,$ (38)
+where $\mathbf{W}[i]\triangleq\mathbf{F}_{\mathrm{RF}}[i]\mathbf{F}_{\mathrm{BB}}\in\mathbb{C}^{N_{\mathrm{t}}\times N_{\mathrm{RF}}}$ is the training combining matrix, $x[i]=1$ represents the transmit pilot, $\mathbf{n}[i]\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{N_{\mathrm{t}}})$ denotes the AWGN.
+The analog combining matrix $\mathbf{F}_{\mathrm{RF}}$ consists of constant-modulus entries randomly drawn from the quantized phase set $\mathcal{F}\triangleq\{\frac{1}{\sqrt{N_{\mathrm{t}}}}e^{\jmath\frac{2\pi b}{2^{B}}}|b=0,1,\ldots,2^{B}-1\}$ and the digital combining matrix is chosen as $\mathbf{F}_{\mathrm{BB}}\triangleq\mathbf{I}_{N_{\mathrm{RF}}}$ for simplicity.
+Collecting all $I$ pilot observations, the complete received pilot sequence at the BS for the $k$-th UE can be formulated as
+$\mathbf{y}_{k}$ $=\mathbf{W}^{H}\mathbf{h}_{\mathrm{RF},k}+\mathbf{n},~\forall k,$ (39)
+where
+$\mathbf{y}_{k}$ $\triangleq[\mathbf{y}_{k}[1]^{T},\mathbf{y}_{k}[2]^{T},\ldots,\mathbf{y}_{k}[I]^{T}]^{T}\in\mathbb{C}^{N_{\mathrm{RF}}I},$ (40)
+$\mathbf{W}$ $\triangleq[\mathbf{W}[1],\mathbf{W}[2],\ldots,\mathbf{W}[I]]\in\mathbb{C}^{N_{\mathrm{t}}\times N_{\mathrm{RF}}I},$
+$\mathbf{n}$ $\triangleq[(\mathbf{W}[1]^{H}\mathbf{n}[1])^{T},(\mathbf{W}[2]^{H}\mathbf{n}[2])^{T},$
+$\quad\quad\quad\quad\quad\quad\ldots,(\mathbf{W}[I]^{H}\mathbf{n}[I])^{T}]^{T}\in\mathbb{C}^{N_{\mathrm{RF}}I}.$
+Considering that the number of RF chains is limited and excessive training overhead must be avoided, the dimension of the received pilot sequence is typically much smaller than the number of BS antennas, i.e., $N_{\mathrm{RF}}I\ll N_{\mathrm{t}}$.
+As a result, directly estimating the high-dimensional RF domain channel $\mathbf{h}_{\mathrm{RF},k}$ from the low dimensional observation $\mathbf{y}_{k}$ becomes highly challenging.
+Fortunately, the scattering cluster channel exhibits inherent sparsity in the angular domain.
+Leveraging this property, we adopt a CS–based channel estimation approach, under which the received signal in (39) can be equivalently represented as
+$\mathbf{y}_{k}$ $=\mathbf{W}^{H}\mathbf{A}_{\mathrm{D}}\mathbf{z}_{k}+\mathbf{n}$ (41)
+$=\mathbf{Q}\mathbf{z}_{k}+\mathbf{n},~\forall k,$
+where $\mathbf{A}_{\mathrm{D}}\in\mathbb{C}^{N_{\mathrm{t}}\times M}$ is the refined dictionary matrix incorporating the effects of pattern-reconfigurable antennas, $\mathbf{z}_{k}\in\mathbb{C}^{M}$ is a sparse angular-domain coefficient vector whose nonzero entries correspond to spatial angles of the rays in the clusters, and $\mathbf{Q}\triangleq\mathbf{W}^{H}\mathbf{A}_{\mathrm{D}}$ is the sensing matrix, which is required to exhibit low mutual coherence and approximate Restricted Isometry Property (RIP) to enable accurate sparse recovery.
+Under such conditions, a CS-based method only requires the total number of effective measurements to exceed the number of propagation paths i.e., $N_{\mathrm{RF}}I\geq L$, to guarantee that $\mathbf{z}_{k}$ can be uniquely identified.
+We subsequently construct a refined dictionary matrix that explicitly embeds the radiation gain of pattern-reconfigurable antennas, yielding a more precise characterization of the spatial-domain channel.
 The conventional dictionary matrix consists of a set of steering vectors corresponding to uniformly distributed candidate angles, given by
-$\mathbf{E}=[\mathbf{e}(\theta_{1}),\mathbf{e}(\theta_{2}),\ldots,\mathbf{e}(\theta_{M})]\in\mathbb{C}^{N_{\mathrm{t}}\times M},$ (39)
-where each column $\mathbf{e}(\theta_{m})$ represents a steering vector pointing toward the angle $\theta_{m}=\pi m/M,~{}m=1,2,\ldots,M$.
-To explicitly account for the influence of the reconfigurable antenna pattern on the channel representation, we refine the dictionary matrix as
-$\mathbf{A}_{\mathrm{D}}(:,m)=\mathbf{F}_{\mathrm{EM}}(:,m)\odot\mathbf{e}(\theta_{m}),~{}\forall m,$ (40)
-where $\mathbf{F}_{\mathrm{EM}}(:,m)$ characterizes the radiation gains of the $N_{\mathrm{t}}$ antennas in the direction of $\theta_{m}$.
+$\mathbf{E}=[\mathbf{e}(\vartheta_{1}),\mathbf{e}(\vartheta_{2}),\ldots,\mathbf{e}(\vartheta_{M})]\in\mathbb{C}^{N_{\mathrm{t}}\times M},$ (42)
+where each column $\mathbf{e}(\vartheta_{m})\in\mathbb{C}^{N_{\mathrm{t}}}$ represents a steering vector pointing toward the angle $\vartheta_{m}=-\frac{\pi}{2}+\frac{\pi m}{M},~m=1,2,\ldots,M$.
+To incorporate the influence of the radiation pattern on the channel representation, we refine the dictionary matrix as
+$\mathbf{A}_{\mathrm{D}}(:,m)=\overline{\mathbf{F}}_{\mathrm{EM}}(:,m)\odot\mathbf{e}(\vartheta_{m}),~\forall m,$ (43)
+where $\overline{\mathbf{F}}_{\mathrm{EM}}\triangleq[\mathbf{f}_{\mathrm{EM},1},\mathbf{f}_{\mathrm{EM},2},\ldots,\mathbf{f}_{\mathrm{EM},N_{\mathrm{t}}}]^{T}\in\mathbb{R}^{N_{\mathrm{t}}\times M}$ stores the radiation patterns of all antennas configured within the current super-frame, and $\overline{\mathbf{F}}_{\mathrm{EM}}(:,m)$ denotes the radiation gains of the $N_{\mathrm{t}}$ antennas in the direction of $\vartheta_{m}$.
 This formulation effectively integrates both radiation pattern effects and angular information, enabling a more accurate estimation of spatial channel characteristics.
+Moreover, to prevent directional bias and ensure uniform scaling across all angular atoms, each column of the refined dictionary is normalized as
+$\mathbf{A}_{\mathrm{D}}(:,m)=\frac{\mathbf{A}_{\mathrm{D}}(:,m)}{\|\mathbf{A}_{\mathrm{D}}(:,m)\|},~\forall m.$ (44)
 Subsequently, the channel estimation problem can be formulated as a sparse recovery problem:
-$\min_{\mathbf{z}_{k}}$ $\ \ \ \|\mathbf{y}_{k}-\mathbf{Q}_{k}\mathbf{z}_{k}\|_{2}$ (41a)
+$\min_{\mathbf{z}_{k}}$ $\ \ \ \|\mathbf{y}_{k}-\mathbf{Q}\mathbf{z}_{k}\|_{2}$ (45a)
 $\mathrm{s.t.}$ $\ \ \ \|\mathbf{z}_{k}\|_{0}=L.$ (45b)
-This problem is typically solved using the orthogonal matching pursuit (OMP) algorithm, which iteratively selects the most relevant dictionary elements to approximate the signal.
+This problem is typically solved using the orthogonal matching pursuit (OMP) algorithm, which iteratively selects the most relevant dictionary elements to approximate the channel.
 The iterations continue until the residual error falls below a predefined threshold or the sparsity constraint is met.
@@
 Thus, the analog beamformer design problem in the medium-timescale is formulated as follows
-$\max_{\mathbf{F}_{\mathrm{RF}}}$ $\ \ \sum^{K}_{k=1}\widetilde{R}_{k}$ (42a)
+$\max_{\mathbf{F}_{\mathrm{RF}}}$ $\ \ \sum^{K}_{k=1}\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{R_{k}\}$ (46a)
 $\mathrm{s.t.}$ $\ \ \mathbf{F}_{\mathrm{RF}}(n,l_{\mathrm{RF}})\in\{\mathcal{F},0\},~\forall n,l_{\mathrm{RF}},$ (46b)
 $\ \ \|\mathbf{F}_{\mathrm{RF}}(n,:)\|_{0}=1,~\forall n,$ (46c)
-where (42a) defines the ergodic sum-rate $\widetilde{R}_{k}\triangleq\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{R_{k}\}$, evaluated over the RF domain channel samples, denoted as $\mathbf{H}_{\mathrm{RF}}\triangleq[\mathbf{h}_{\mathrm{RF},1},\mathbf{h}_{\mathrm{RF},2},\ldots,\mathbf{h}_{\mathrm{RF},K}]\in\mathbb{C}^{N_{\mathrm{t}}\times K}$ with $\mathbf{h}_{\mathrm{RF},k}$ given in (37).
+where the objective function (46a) defines the ergodic sum-rate evaluated over the RF domain channel samples, denoted as $\mathbf{H}_{\mathrm{RF}}\triangleq[\mathbf{h}_{\mathrm{RF},1},\mathbf{h}_{\mathrm{RF},2},\ldots,\mathbf{h}_{\mathrm{RF},K}]\in\mathbb{C}^{N_{\mathrm{t}}\times K}$ with $\mathbf{h}_{\mathrm{RF},k}$ given in (37).
 Notably, in this RF domain analog beamformer design problem, the power constraint is not explicitly imposed since the digital beamformer can be subsequently adjusted to satisfy it.
@@
 $\mathbf{F}_{\mathrm{RF}}\triangleq\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}},$ (47)
-where $\mathbf{S}_{\mathrm{RF}}\in\{0,1\}^{N_{\mathrm{t}}\times N_{\mathrm{RF}}2^{B}}$ is the RF domain phase-shift selection matrix and $\mathbf{F}_{\mathrm{set}}$ is defined as $\mathbf{F}_{\mathrm{set}}\triangleq\mathbf{I}_{N_{\mathrm{RF}}}\otimes\mathbf{f}_{\mathrm{set}}$ with
-$\mathbf{f}_{\mathrm{set}}\triangleq\frac{1}{\sqrt{N_{\mathrm{t}}}}\left[1,e^{-\jmath\frac{2\pi}{2^{B}}},\ldots,e^{-\jmath\frac{2\pi(2^{B}-1)}{2^{B}}}\right]^{T},$ (44)
-containing all possible phase-shifts for a low-resolution PS.
+where $\mathbf{S}_{\mathrm{RF}}\in\{0,1\}^{N_{\mathrm{t}}\times N_{\mathrm{RF}}2^{B}}$ is the RF domain phase-shift selection matrix and $\mathbf{F}_{\mathrm{set}}$ is defined as $\mathbf{F}_{\mathrm{set}}\triangleq\mathbf{I}_{N_{\mathrm{RF}}}\otimes\mathbf{f}_{\mathrm{set}}$ with $\mathbf{f}_{\mathrm{set}}$ containing all possible quantized phase-shifts, yielding
+$\mathbf{f}_{\mathrm{set}}\triangleq\frac{1}{\sqrt{N_{\mathrm{t}}}}\left[1,e^{-\jmath\frac{2\pi}{2^{B}}},\ldots,e^{-\jmath\frac{2\pi(2^{B}-1)}{2^{B}}}\right]^{T}.$ (48)
 The binary matrix $\mathbf{S}_{\mathrm{RF}}$ encodes phase-shift selections, where $\mathbf{S}_{\mathrm{RF}}(n,l_{\mathrm{RFB}})=1$ indicates that the $n$-th antenna is assigned to the $l_{\mathrm{RFB}}/2^{B}$-th RF chain with the corresponding phase-shift value given by $\mathbf{F}_{\mathrm{set}}(l_{\mathrm{RFB}},l_{\mathrm{RFB}}/2^{B})$.
 With $\mathbf{F}_{\mathrm{set}}$ predefined, the analog beamformer design problem (46) can be equivalently formulated as
-$\max_{\mathbf{S}_{\mathrm{RF}}}$ $\ \ \ \sum^{K}_{k=1}\widetilde{R}_{k}$ (45a)
+$\max_{\mathbf{S}_{\mathrm{RF}}}$ $\ \ \ \sum^{K}_{k=1}\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{R_{k}\}$ (49a)
 $\mathrm{s.t.}$ $\ \ \ \mathbf{S}_{\mathrm{RF}}(n,l_{\mathrm{RFB}})\in\{0,1\},~\forall n,l_{\mathrm{RFB}},$ (49b)
@@
 where the sum-rate expression in (49a) can be recast as
-$\widetilde{R}_{k}=\log(1+\frac{|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},k}|^{2}}{\sum_{j=1,j\neq k}^{K}|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},j}|^{2}+\sigma_{k}^{2}}),~{}\forall k.\$ (46)
+$R_{k}\!\!=\!\log_{2}\!\left(\!1\!\!+\!\frac{|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},k}|^{2}}{\sum_{j=1,j\neq k}^{K}\!|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},j}|^{2}\!\!+\!\!\sigma_{k}^{2}}\!\right)\!,\forall k.\!\!\!\!\\$ (50)
 The primary challenges in solving this problem stem from the stochastic nature of the objective function (49a), which lacks a closed-form expression, and the non-smooth, non-convex Boolean constraint (49b).
-To address these issues, we employ a stochastic successive convex approximation (SSCA) framework [25], which iteratively refines the beamformer based on observed channel realizations.
+To address these issues, we employ a stochastic successive convex approximation (SSCA) framework [28], which recursively refines the beamformer based on observed channel realizations.
 The expectation in the objective function is approximated using a quadratic surrogate function, allowing for online adaptation without requiring explicit statistical knowledge of the channel.
@@
 First, we construct a recursive quadratic surrogate function for the objective function (49a), leveraging both current and past channel samples.
-This iterative approach enables the BS to gradually learn the statistical properties of the channel by continuously accumulating channel observations, thereby enhancing beamforming performance over time and ensuring convergence.
+This recursive approach enables the BS to gradually learn the statistical properties of the channel by continuously accumulating channel observations, thereby enhancing beamforming performance over time and ensuring convergence.
 For convenience in the following expressions, we define $g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\triangleq R_{k}$, and introduce $\mathbf{s}_{\mathrm{RF}}\triangleq\mathrm{vec}(\mathbf{S}_{\mathrm{RF}})$.
-At the $t_{\mathrm{M}}$-th recursion, based on the acquired channel sample $\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$ and the previous analog beamformer $\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$, we can substitute (45a) with a quadratic surrogate function as follows
-$f^{(t_{\mathrm{M}})}(\mathbf{s}_{\mathrm{RF}})\!=\!v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}+(\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})})^{T}(\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})})-\tau\|\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}\|^{2},$ (47)
-where $\tau>0$ is a constant ensuring strong convexity of $\tau\|\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}\|^{2}$.
+At the $t_{\mathrm{M}}$-th recursion, based on the acquired channel sample $\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$ and the previous RF domain phase-shift selection matrix $\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)}$, we can substitute (49a) with a quadratic surrogate function as follows
+$f^{(t_{\mathrm{M}})}\!(\mathbf{s}_{\mathrm{RF}}\!)\!=\!v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}+(\!\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}\!)^{T}\!(\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}\!-\!1)})-\tau\|\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}\!-\!1)}\!\|^{2},$ (51)
+where $\tau>0$ is a constant ensuring strong convexity of $\tau\|\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)}\|^{2}$.
 The term $v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}$ serves as an approximation of the objective function value $\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\}$, which is updated recursively as
-$v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}=(1-\eta^{(t_{\mathrm{M}})})v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}}-1)}+\eta^{(t_{\mathrm{M}})}g_{0}(\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})};\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}),$ (48)
-where $\eta^{(t_{\mathrm{M}})}\in(0,1]$ is a properly chosen sequence to satisfy convergence requirements.
-Similarly, the gradient term $\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}$ is an approximation of the partial derivative of $\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\}$ with respect to $\mathbf{s}_{\mathrm{RF}}$, evaluated at $\mathbf{s}_{\mathrm{RF}}=\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$.
+$v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}=(1-\eta^{(t_{\mathrm{M}})})v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}}-1)}+\eta^{(t_{\mathrm{M}})}g_{0}(\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)};\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}),$ (52)
+where $\eta^{(t_{\mathrm{M}})}$ is a properly chosen sequence to satisfy convergence requirements [28].
+Similarly, the gradient term $\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}$ is an approximation of the partial derivative of $\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\}$ with respect to $\mathbf{s}_{\mathrm{RF}}$, evaluated at $\mathbf{s}_{\mathrm{RF}}=\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)}$.
 This is recursively computed as
-$\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}=(1-\eta^{(t_{\mathrm{M}})})\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}}-1)}+\eta^{(t_{\mathrm{M}})}\nabla_{\mathbf{s}_{\mathrm{RF}}}g_{0}(\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})};\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}).$ (49)
+$\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}\!\!=\!\!(1\!-\!\eta^{(t_{\mathrm{M}})})\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}}\!-\!1)}\!+\!\eta^{(t_{\mathrm{M}})}\nabla_{\mathbf{s}_{\mathrm{RF}}}g_{0}(\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}\!-\!1)};\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}).$ (53)
 The gradient of $g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})$ with respect to $\mathbf{s}_{\mathrm{RF}}$ is given by
-$\nabla_{\mathbf{s}_{\mathrm{RF}}}g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})=\sum_{k=1}^{K}\left(\frac{\sum_{i}^{K}\mathbf{e}_{k,i}}{\mit{\Gamma_{k}}}-\frac{\sum_{i\neq k}^{K}\mathbf{e}_{k,i}}{\mit{\Gamma_{-k}}}\right),$ (50)
+$\nabla_{\mathbf{s}_{\mathrm{RF}}}g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})=\sum_{k=1}^{K}\left(\frac{\sum_{i}^{K}\mathbf{e}_{k,i}}{{\Gamma_{k}}}-\frac{\sum_{i\neq k}^{K}\mathbf{e}_{k,i}}{{\Gamma_{-k}}}\right),$ (54)
 where we define
-$\mit{\Gamma_{k}}\!$ $=\!\sum_{i}^{K}\left|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},i}\right|^{2}+\sigma_{k}^{2},~{}\forall k,$ (51a)
-$\mit{\Gamma_{-k}}\!$ $=\!\sum_{i\neq k}^{K}\left|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},i}\right|^{2}+\sigma_{k}^{2},~{}\forall k,$ (51b)
+${\Gamma_{k}}\!$ $=\!\sum_{i}^{K}\left|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},i}\right|^{2}+\sigma_{k}^{2},~\forall k,$ (55a)
+${\Gamma_{-k}}\!$ $=\!\sum_{i\neq k}^{K}\left|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},i}\right|^{2}+\sigma_{k}^{2},~\forall k,$ (55b)
 $\mathbf{e}_{k,i}\!$ $=2\Re\{\mathrm{vec}((\mathbf{h}_{\mathrm{RF},k}\mathbf{h}_{\mathrm{RF},k}^{H})\mathbf{S}_{\mathrm{RF}}(\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},i}\mathbf{f}_{\mathrm{BB},i}^{H}\mathbf{F}_{\mathrm{set}}^{H}))\}.$ (55c)
@@
 To deal with the Boolean constraint (49b), we employ the same technique as specified in (33)-(35). As a result, the analog beamformer design problem is formulated as
-$\max_{\mathbf{S}_{\mathrm{RF}}}$ $\ \ \ f^{(t_{\mathrm{M}})}(\mathbf{s}_{\mathrm{RF}})+\varrho_{2}(2\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}-\mathbf{1})^{T}\mathbf{s}_{\mathrm{RF}}$ (52a)
+$\max_{\mathbf{S}_{\mathrm{RF}}}$ $\ \ \ f^{(t_{\mathrm{M}})}(\mathbf{s}_{\mathrm{RF}})+\varrho_{2}(2\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)}-\mathbf{1})^{T}\mathbf{s}_{\mathrm{RF}}$ (56a)
 $\mathrm{s.t.}$ $\ \ \ \mathbf{S}_{\mathrm{RF}}(n,l_{\mathrm{RFB}})\in[0,1],~\forall n,l_{\mathrm{RFB}},$ (56b)
@@
 This update is performed using the following recursive expression:
-$\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}=(1-\eta^{(t_{\mathrm{M}})})\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}+\eta^{(t_{\mathrm{M}})}\mathbf{\overline{S}}^{(t_{\mathrm{M}})}_{\mathrm{RF}}.$ (53)
-With the updated phase-shift selection matrix $\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}$, the analog beamformer is then constructed as $\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}=\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}\mathbf{F}_{\mathrm{set}}$.
+$\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}=(1-\eta^{(t_{\mathrm{M}})})\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)}+\eta^{(t_{\mathrm{M}})}\mathbf{\overline{S}}^{(t_{\mathrm{M}})}_{\mathrm{RF}}$.
+(57)
+With the updated phase-shift selection matrix $\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$, the analog beamformer is then constructed as $\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}})}=\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}\mathbf{F}_{\mathrm{set}}$.
 III-D Short-timescale Digital Beamformer Design
@@
 The resulting digital beamformer is given by
-$\mathbf{\widetilde{F}}_{\mathrm{BB}}^{(t_{\mathrm{S}})}=(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})})^{H}\left(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})}(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})})^{H}+\sigma_{k}^{2}\mathbf{I}_{K}\right)^{-1}.$ (54)
+$\mathbf{\widetilde{F}}_{\mathrm{BB}}^{(t_{\mathrm{S}})}=(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})})^{H}\left(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})}(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})})^{H}+\bm{\Lambda}\right)^{-1},$ (58)
+where $\bm{\Lambda}\triangleq\mathrm{diag}\{\sigma_{1}^{2},\sigma_{2}^{2},\ldots,\sigma_{K}^{2}\}$.
 Finally, to ensure compliance with the total power constraint, the digital beamformer is normalized as
@@
     • Calculate the channel covariance matrix $\mathbf{R}_{k}^{(t_{\mathrm{L}})},\forall k$ by (19).
-    • Derive the APS $\rho^{(t_{\mathrm{L}})}_{k}(\theta),\forall k$ as (19).
+    • Derive the APS $\rho^{(t_{\mathrm{L}})}_{k}(\vartheta_{m}),\forall k$ as (20).
     • Estimate the nominal angles $\theta_{k,c},\forall k,c,$ and average channel gains $\alpha_{k,c},\forall k,c,$ of the scattering clusters.
@@
   • Analog beamformer design
-    • Update the surrogate function (47) based on $\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})},\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$.
+    • Update the surrogate function (51) based on $\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})},\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)}$.
     • Calculate the optimal $\mathbf{\overline{S}}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$ using (56).
-    • Update $\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}$ according to (53).
-    • Construct the analog beamformer $\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}$ by (43).
+    • Update $\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$ according to (57).
+    • Construct the analog beamformer $\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$ by (47).
  Step 3: Short-timescale BB domain design at the $t_{\mathrm{S}}$-th time-slot
@@
 Based on the above derivations, the proposed tri-timescale tri-hybrid beamforming design is straightforward and summarized in Algorithm 1.
-After appropriately initializing the beamforming matrices $\mathbf{F}_{\mathrm{EM}}$, $\mathbf{F}_{\mathrm{RF}}$ and $\mathbf{F}_{\mathrm{BB}}$ across the three domains, the algorithm iteratively acquires the corresponding channel information and updates these matrices over the three timescales, respectively.
+After appropriately initializing the beamforming matrices $\mathbf{F}_{\mathrm{EM}}$, $\mathbf{F}_{\mathrm{RF}}$ and $\mathbf{F}_{\mathrm{BB}}$ across the three domains, the algorithm alternately acquires the corresponding channel information and updates these matrices over the three timescales, respectively.
 III-E.2 Complexity Analysis
@@
 The primary computational complexity lies in calculating the EM domain pattern selection matrix $\mathbf{S}_{\mathrm{EM}}$ in (36) over the long-timescale, and determining the phase-shift selection matrix $\mathbf{S}_{\mathrm{RF}}$ in (56) over the medium-timescale.
-Specifically, the computational complexity of solving problem (36) in each super-frame over the long-timescale, which involves an $N_{\mathrm{t}}P$-dimensional variable and requires handling $2N_{\mathrm{t}}P$ second-order cone (SOC) constraints, is of order $\mathcal{O}\{\sqrt{2N_{\mathrm{t}}P+N_{\mathrm{t}}+1}N_{\mathrm{t}}P(N_{\mathrm{t}}^{2}P^{2}+2N_{\mathrm{t}}P+N_{\mathrm{t}})\}$ [34].
-Similarly, the computational complexity of solving problem (52) in each frame during the medium-timescale, which entails calculating an $N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}$-dimensional variable along with $2N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}$ SOC constraints, is of order $\mathcal{O}\{\sqrt{2N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}+N_{\mathrm{t}}+1}N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}(N_{\mathrm{t}}^{2}N_{\mathrm{RF}}^{2}2^{2B}+2N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}+N_{\mathrm{t}})\}$.
+Specifically, the computational complexity of solving problem (36) in each super-frame over the long-timescale with an $N_{\mathrm{t}}P$-dimensional variable, is of order $\mathcal{O}\{N_{\mathrm{t}}^{3.5}P^{3.5}\}$ [38].
+Similarly, the computational complexity of solving problem (56) in each frame during the medium-timescale with an $N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}$-dimensional variable is of order $\mathcal{O}\{N_{\mathrm{t}}^{3.5}N_{\mathrm{RF}}^{3.5}2^{3.5B}\}$.
 III-E.3 Pilot Overhead Analysis
 The pilot overhead in the proposed scheme primarily stems from channel sampling at the medium-timescale and short-timescale stages.
-In the medium-timescale, CS-based channel estimation requires $N_{\mathrm{Beam}}K/N_{\mathrm{RF}}$ pilot symbols to obtain the RF domain high-dimensional statistical channel sample $\mathbf{h}_{\mathrm{RF},k}^{(t_{\mathrm{M}})},\forall k$, as described in Section III-C.In the short-timescale, $N_{\mathrm{RF}}K$ pilot symbols are needed for estimating the BB domain low-dimensional effective channel $\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})}$, as described in Section III-D.
-Therefore, the total pilot overhead across the channel statistics coherence time is expressed as $(N_{\mathrm{Beam}}K/N_{\mathrm{RF}}+N_{\mathrm{RF}}KT_{\mathrm{S}})T_{\mathrm{M}}T_{\mathrm{L}}$.
-This is notably less than that of the real-timescale beamforming approaches, which requires $(N_{\mathrm{Beam}}K/N_{\mathrm{RF}}+N_{\mathrm{RF}}K)T_{\mathrm{S}}T_{\mathrm{M}}T_{\mathrm{L}}$.
+In the medium-timescale, CS-based channel estimation requires $I$ pilot symbols $(N_{\mathrm{RF}}I\gg L)$ for each UE to obtain the RF domain high-dimensional statistical channel sample $\mathbf{h}_{\mathrm{RF},k}^{(t_{\mathrm{M}})}$, as described in Section III-C.
+In the short-timescale, $N_{\mathrm{RF}}K$ pilot symbols are needed for estimating the BB domain low-dimensional effective channel $\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})}$, as described in Section III-D.
+Therefore, the total pilot overhead across the channel statistics coherence time is expressed as $(KI+KT_{\mathrm{S}})T_{\mathrm{M}}T_{\mathrm{L}}$.
+This is notably less than that of the real-timescale beamforming approaches, which requires $(KI+K)T_{\mathrm{S}}T_{\mathrm{M}}T_{\mathrm{L}}$.
+III-E.4 Initialization
+During the initialization phase, we set all antenna patterns to point toward the boresight direction for the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$, adopt a random phase initialization for the analog beamformer $\mathbf{F}_{\mathrm{RF}}$, and employ a normalized identity matrix for the digital beamformer $\mathbf{F}_{\mathrm{BB}}$.
+These initial settings serve as a neutral starting point for the subsequent optimization.
 IV Simulation Results
@@
 The low-resolution PS used in the RF domain has quantized phase controlled by $B=3$ bits.
-Without loss of generality, we assume that the number of UEs $K$ is equal to the number of RF chains $N_{\mathrm{RF}}$.
-In the EM domain, each pixel-based reconfigurable antenna enables the generation of $P=7$ distinct directional radiation patterns by adjusting the electronic switches between pixels [8], with $M=180$ uniform sampling points in the azimuth dimension.
+For simplicity and to focus on the tri-timescale design, we set $N_{\mathrm{RF}}=K$ in simulations and the framework applies to $N_{\mathrm{RF}}\geq K$ as well.
+In the EM domain, each pixel-based reconfigurable antenna enables the generation of $P=7$ distinct directional radiation patterns by adjusting the electronic switches between pixels [12], with $M=180$ uniform sampling points in the azimuth dimension [^3] The prototype reconfigurable antenna in [12] currently has an aperture exceeding the conventional half-wavelength spacing, introducing practical challenges such as mutual coupling and grating lobes, which we leave for future investigation.
+In this study, we adopt half-wavelength spacing to facilitate fair comparisons with conventional array designs.
+With ongoing advancements in reconfigurable antenna technology, recent studies and prototypes have convincingly demonstrated the feasibility of smaller apertures compatible with half-wavelength spacing, thereby addressing these practical concerns..
 In this scenario, there are $K=2$ UEs, all operating under a scattering cluster channel with $C=2$ clusters, each of which consists of $L_{c}=3$ multi-path rays.
-The nominal AoDs $\overline{\theta}_{k,c}$ of these scattering clusters are randomly distributed within $[-\pi/3,\pi/3]$, while the angle spread is set to $\varsigma_{k,c}=\pi/36$.
-The path loss is expressed as $\alpha_{k,c,l_{c}}=10^{-\frac{C_{0}}{10}}(\frac{r_{k,c,l_{c}}}{D_{0}})^{-\kappa_{k,c,l_{c}}}$ with $\kappa_{k,c,l_{c}}\in[2.5,3]$ being the attenuation factor, $C_{0}=30$ dB, $D_{0}=1$m.
-The propagation distance $r_{k,c,l_{c}}$ is composed of the nominal distance $\overline{r}_{k,c}\in[50\mathrm{m},100\mathrm{m}]$ and $\iota=0.5\mathrm{m}$.
+The nominal AoDs $\overline{\theta}_{k,c}$ of these scattering clusters are randomly distributed within $\pi/2,\pi/2]$, while the angle spread is set to $\varsigma_{k,c}=\pi/36$.
+The large-scale path loss is expressed as $\alpha_{k,c}=10^{-\frac{C_{0}}{10}}(\frac{r_{k,c}}{D_{0}})^{-\kappa_{k,c}}$ with $\kappa_{k,c}\in[2.5,3]$ being the attenuation factor, $C_{0}=30$ dB, $D_{0}=1$m.
+The propagation distance $r_{k,c}$ of each scattering cluster is randomly distributed in $r_{k,c}\in[50\mathrm{m},100\mathrm{m}]$.
 The noise power of UEs is set as $\sigma_{k}^{2}=-70$dBm.
@@
 (b) Medium-timescale.
-Figure 5: Convergence performance of the long-timescale radiation beamformer design and medium-timescale analog beamformer design.
+Figure 5: Convergence performance of the radiation beamformer design and analog beamformer design.
+Figure 6: The sum-rate versus the transmit power $P_{\mathrm{t}}$.
 The convergence performance of the proposed tri-timescale tri-hybrid beamforming design is illustrated in Fig.
@@
 5 demonstrate that both algorithms exhibit satisfactory convergence speed, indicating an efficient level of computational complexity.
-Figure 6: The sum-rate versus the transmit power $P_{\mathrm{t}}$.
+Figure 7: The sum-rate versus the distance and angle of UE.
+Figure 8: The sum-rate and channel estimation overhead versus the number of frames $T_{\mathrm{M}}$.
 Fig.
-6 illustrates the sum-rate performance versus transmit power for systems employing reconfigurable and conventional antennas.
-Specifically, the proposed tri-hybrid beamforming architecture with reconfigurable antennas and a dynamic-subarray scheme is denoted as “RA, DS”.
-For comparison, hybrid beamforming using conventional antennas with a dynamic-subarray is labeled “CA, DS”, and the fixed-subarray architectures with reconfigurable and conventional antennas are denoted as “RA, FS” and “CA, FS”, respectively.
-Performance is evaluated for systems with $N_{\mathrm{t}}=32$ and $N_{\mathrm{t}}=16$ antennas.
-As shown in Fig.
-6, the proposed “RA, DS” algorithm consistently outperforms its competitors.
-In particular, reconfigurable antennas yield a power gain of approximately 7 dB compared to conventional antennas, while the dynamic-subarray scheme provides an additional gain of about 2 dB over the fixed-subarray approach.
-These results confirm that the tri-hybrid beamforming architecture achieves superior communication performance by virtue of its enhanced flexibility in radiation pattern selection in the EM domain and subarray selection in the RF domain.
-Figure 7: The sum-rate and channel estimation overhead versus the number of frames $T_{\mathrm{M}}$.
+6 illustrates the sum-rate performance versus transmit power for systems employing reconfigurable antennas (RA) and conventional antennas (CA).
+The proposed tri-hybrid beamforming architecture with reconfigurable antennas and dynamic-subarray selection is labeled as “RA, DS”.
+For benchmarking purposes, fully-digital beamforming with reconfigurable antennas denoted as “RA, FD” serves as the upper performance bound and corresponds to a real-timescale optimization, while hybrid beamforming with fixed-subarray selection denoted as“RA, FS” represents the lower bound.
+Specifically, the “RA, DS and “RA, FS” configurations implement the proposed tri-timescale optimization framework, where EM domain, RF domain, and BB domain beamformers are optimized across distinct timescales.
+For comparison, we also include three schemes employing conventional antennas: “CA, DS”, “CA, FD”, and “CA, FS”.
+Among them, “CA, FD” corresponds to a real-timescale design, whereas “CA, DS” and “CA, FS” employ a conventional two-timescale optimization approach, optimizing analog beamformers over a long timescale and digital beamformers over a short timescale [28]-[30].
+As observed in Fig.
+6, the proposed “RA, DS” approach consistently outperforms other schemes.
+In particular, the use of reconfigurable antennas achieves approximately a 7 dB gain compared to conventional antennas.
+Furthermore, employing dynamic-subarray selection yields an additional gain of roughly 2 dB over fixed-subarray schemes.
+These results validate that the proposed tri-hybrid beamforming architecture, benefiting from adaptive EM domain radiation patterns and flexible RF domain subarray selections, significantly enhances overall communication performance.
+(a) Long-timescale.
+(b) Medium-timescale.
+Figure 9: The performance evaluation of channel estimation in the long-timescale and medium-timescale.
+Figure 10: The sum-rate versus the number of angular sampling points $M$.
 For the proposed tri-timescale tri-hybrid beamforming design, increasing the channel sampling rate in the medium-timescale can enhance sum-rate performance by enabling more frequent updates of channel statistics for multi-path rays and scattering cluster cores.
@@
 To quantitatively evaluate the impact of the channel sampling rate on both system performance and pilot overhead, we gradually increase the number of frames per super-frame in the medium-timescale, effectively increasing the channel sampling rate, as illustrated in Fig.
-7.
+8.
 By analyzing the pilot overhead expression derived in Section III-E and observing the trend in Fig.
-7, we conclude that the proposed tri-timescale scheme significantly reduces pilot overhead compared to the real-time scheme, which requires high-dimensional CSI estimation in every time-slot.
+8, we conclude that the proposed tri-timescale scheme significantly reduces pilot overhead compared to the real-time scheme, which requires high-dimensional CSI estimation in every time-slot.
 More importantly, as the number of frames increases, the sum-rate performance of the tri-timescale beamforming scheme steadily improves and eventually saturates at a satisfactory level, albeit with a monotonic increase in channel estimation overhead.
-In particular, setting the number of frames to $T_{\mathrm{M}}=50$ achieves a favorable balance between performance and overhead, a configuration adopted in the simulations.
-Figure 8: The impact of channel estimation configurations on sum-rate performance.
-The channel estimation algorithms in both the EM and RF domains are critical components of the proposed tri-timescale framework.
-Fig.
-8 illustrates the impact of different channel estimation configurations on sum-rate performance under various system settings.
-The results indicate that, regardless of the number of propagation paths $L_{\mathrm{c}}$, the sum-rate achieved with estimated CSI is highly comparable to that with perfect CSI.
-This finding underscores the accuracy and robustness of our channel estimation algorithms within the proposed framework, even under complex propagation conditions.
-Moreover, even with a reduced angular resolution $N_{\mathrm{Beam}}$, the channel characteristics are reliably captured, and acceptable sum-rate performance is achieved, further validating the effectiveness and reliability of the proposed approach.
-Figure 9: The sum-rate versus the distance and angle of UE.
-Finally, we validate the effectiveness of the proposed scheme by evaluating the impact of user location on the overall performance of the tri-hybrid beamforming with reconfigurable antennas (RA) and the hybrid beamforming with conventional antennas (CA).
+In particular, selecting the number of frames as $T_{\mathrm{M}}=50$ achieves an optimal balance between performance and overhead, and this configuration is therefore employed in our simulations.
+Figure 11: The sum-rate versus the number of antennas and UEs.
+Next, we validate the effectiveness of the proposed scheme by evaluating the impact of user location on the overall performance of the tri-hybrid beamforming with reconfigurable antennas and the hybrid beamforming with conventional antennas.
 Specifically, Fig.
@@
 These results underscore the robustness and versatility of the proposed tri-hybrid beamforming approach, demonstrating that dynamic pattern reconfiguration not only compensates for variations in user location but also substantially enhances overall system performance and reliability across a wide range of deployment scenarios.
+The channel estimation algorithms in both the EM and RF domains are critical components of the proposed tri-timescale framework.
+As shown in Fig.
+9 (a), the normalized APS obtained by the Capon method is depicted as a function of the angle.
+Two dominant peaks can be clearly observed, corresponding to the scattering clusters in the angular domain.
+The estimated nominal angles, marked by red dots, are well aligned with the peaks of the Capon APS, indicating accurate cluster angle estimation in the long timescale.
+The normalized mean square error (NMSE) performance of channel estimation is evaluated versus SNR for different schemes.
+The CS-based channel estimation using dictionary with pattern information consistently outperforms its counterpart without pattern information, demonstrating the benefit of exploiting pattern-aware dictionaries in the medium-timescale channel estimation process.
+In Fig.
+10, we investigate how the number of angular sampling points $M$ affects the achievable sum-rate.
+It is evident that the system sum-rate monotonically increases with $M$, since a finer angular grid provides a more accurate discretization of continuous radiation patterns, thus improving the matching precision between radiation gains and user angles.
+Nevertheless, the incremental benefit gradually becomes negligible as $M$ increases beyond a certain point (around $M=60$ for perfect CSI and $M=180$ for estimated CSI), indicating limited value in excessively dense angular sampling.
+Specifically, after comprehensive consideration, the choice of $M=180$, adopted in earlier simulations, achieves a favorable balance between performance improvement and computational complexity, effectively capturing most performance gains without incurring excessive complexity.
+To further examine the scalability of the proposed tri-hybrid beamforming architecture, Fig.
+11 illustrates the achievable sum-rate performance under varying numbers of BS antennas and UEs.
+The results in Fig.
+11 (a) clearly underscore the hardware efficiency and substantial array-size reductions achievable with reconfigurable antennas, highlighting their considerable potential for supporting energy-efficient next-generation wireless systems without compromising communication performance.
+Additionally, as depicted in Fig.
+11 (b), the proposed tri-hybrid beamforming consistently outperforms both baseline hybrid beamforming and conventional antenna schemes across all considered user loads, demonstrating robust scalability.
+These comprehensive simulation results thus strongly validate the practicality and scalability of the proposed tri-hybrid architecture for future large-scale multi-user wireless deployments.
 V Conclusions
@@
 This hierarchical design effectively balances system performance with practical constraints, such as channel estimation overhead and computational complexity.
-Extensive simulation results validate the superiority of the proposed scheme over conventional hybrid beamforming architectures, highlighting its potential in future wireless communication systems.
+Numerical results confirmed that the proposed design achieves around 7 dB sum-rate improvement, while reducing pilot overhead by nearly an order of magnitude compared with real-time joint beamforming schemes.
+The demonstrated scalability and performance gains underscore the practical feasibility and effectiveness of our approach for future large-scale wireless systems.
+Future work will address practical hardware aspects and extend this framework to wideband and near-field scenarios.
+Future research directions include addressing practical hardware challenges associated with reconfigurable antennas, such as severe mutual coupling effects resulting from half-wavelength spacing.
+Incorporating multi-port circuit theory to model and mitigate mutual coupling through joint radiation pattern and hybrid beamforming design is a promising pathway.
+Furthermore, larger antenna spacing can introduce strong grating lobes in practice.
+To address this issue, it will be essential to explore sparse or non-uniform array configurations, irregular sampling methods, and integrated radiation pattern optimization schemes.
+Investigating these aspects will further extend the practical applicability and deepen the theoretical understanding of tri-hybrid beamforming architectures for realistic deployment scenarios.